### v2

# In Search of the Long-Tail: Systematic Generation of Long-Tail Inferential Knowledge via Logical Rule Guided Search

Abstract
State-of-the-art LLMs outperform humans on reasoning tasks such as Natural Language Inference. Recent works evaluating LLMs note a marked performance drop on input data from the low-probability distribution, i.e., the long-tail. Therefore, we focus on systematically generating statements involving long-tail inferential knowledge for more effective evaluation of LLMs in the reasoning space. We first propose a novel framework Logic-Induced-Knowledge-Search (LINK[Uncaptioned image]) [^1] https://github.com/INK-USC/LINK/ that generates factually correct and long-tail knowledge statements grounded on symbolic rule templates; LINK effectively generates data in the long-tail distribution that zero-shot prompted LLMs are unable to reach, and outperforms zero-shot GPT4 on factual correctness by 5%. We further use the data generated by LINK to construct a dataset Logic-Induced-Long-Tail (LINT) [^2] https://doi.org/10.5281/zenodo.10691770 that can be used to evaluate downstream models on the long-tail distribution; LINT contains 108K knowledge statements spanning four domains. We use LINT to test LLMs on an entailment classification task and find that model performances drop by as high as 5% in the long-tail distribution compared to head distribution. Our work shows the utility of evaluating models in the long-tail distribution, and calls for more research on generating evaluation data in the long-tail distribution.
[…]
In Search of the Long-Tail: Systematic Generation of Long-Tail Inferential Knowledge via Logical Rule Guided Search
[…]
Figure 1: An illustration of the distribution of statements generated by different methods, after being filtered on factual correctness by human evaluation. For the same symbolic rule, all methods generate the same number of statements, but LINK(Refer to caption) has more correct statements than zero-shot ChatGPT and zero-shot GPT4(Refer to caption).  LINK long-tail generations clearly lie in a more long-tail part of the probability distribution than the zero-shot GPT4 generations.
1 Introduction
Large language models (LLMs) have approached broadly human-level performance on an increasing number of benchmarks (OpenAI, 2023; Ouyang et al., 2022). The problem of evaluating models that potentially outperform (non-expert) humans on most skills relevant to the task seems more imminent than ever (Bowman et al., 2022). Natural Language Inference (Bowman et al., 2015; Zellers et al., 2019), which requires extensive knowledge and complex reasoning about entities and events, is one of such tasks on which LLMs have impressive performance (Achiam et al., 2023; Touvron et al., 2023b; Jiang et al., 2023). Recent works on evaluating LLMs on challenge reasoning sets notice a marked decline in performance when facing input data from the “long-tail”, i.e., the low-probability part of the natural language’s distribution [^3] See Section 2.1 for discussion on the relationship between the probability distribution of natural language and LLMs.(McCoy et al., 2023; Razeghi et al., 2022).
To have a well-rounded evaluation of LLMs within the reasoning space, one needs access to inferential knowledge statements (Figure 1) collected systematically from the long-tail to form challenging evaluation sets (i.e., long-tail inferential knowledge). However, so far there has not been any systematic framework that targets generating data in the long-tail of probability distribution of language models. Crowdsourcing long-tail data is difficult because of human cognitive bias (Tversky and Kahneman, 1973, 1974), and LLMs’ generation in the long-tail distribution is hindered by their pretraining task of “most likely" next token (McCoy et al., 2023).
In this paper, we propose a logical rule-guided long-tail knowledge generation framework, Logic-Induced-Knowledge-Search (LINK[Uncaptioned image]), for generating inferential knowledge statements that are both factually correct and in the long-tail  – a task that is fundamentally challenging to SoTA LLMs. To achieve the two goals simultaneously, we first ground the generation of inferential knowledge statements with symbolic rules, which consist of variables and predicates that specify the relationship between variables. The knowledge statements are ensured to be valid as long as the filled values of the variables satisfy all predicates in the rule. Doing so significantly simplifies the task from sentence generation to value search with the given rules. Specifically, we propose knowledge beam search (Figure 2) to utilize the modularity of symbolic rules by searching for values of the variables one by one in a designated order – an algorithm inspired by sequential beam search. At each step, we prompt an LLM for plausible values that satisfy all predicates in which the variable appears. The step-wise search design also allows for granular control of the statement’s correctness and likelihood as it incorporates a critic model that checks for data type conformity and factual correctness and a reranker model that ensures that the values fall into the long-tail of the distribution.
LINK shows clear advantage in generating long-tail knowledge statements compared to prompting LLMs (Figure 1). We find that even with symbolic rules presented in the prompt, e.g., by instructing LLM to fill in values for the rules, the prompt-based methods using ChatGPT (gpt-3.5-turbo) and GPT4 (gpt-4) still struggle to generate knowledge statements that are both correct and in the long-tail distribution. This implies that the prompt-based methods may suffer from directly retrieving long-tail knowledge from the natural language distribution. In addition, human annotators find that compared to LINK, the statements generated by prompt-based methods are prone to have more factual errors. Further analysis shows (1) the key contribution to the long-tail generation quality comes from the LINK pipeline, (2) LINK can be effectively and flexibly applied across different knowledge models, and (3) post-hoc reranking does not reach into the long-tail distribution as LINK.
Using LINK, we curate a dataset, Logic-Induced-Long-Tail (LINT), which contains 108k knowledge statements spanning across four different domains, presented in a premise, conclusion format (Table 1). LINT can be a useful resource for the community to more thoroughly test LLMs’ capabilities in the long-tail distribution. We use a subset of LINT to create an evaluation task in the form of entailment classification, where LLMs are asked to judge true or false on a knowledge statement. In the long-tail test set, GPT4’s capability in identifying incorrect knowledge drop by as high as 3% with zero-shot prompting, and as high as 4% with Chain-of-Thought prompting (Wei et al., 2022). GPT4’s performance gap from human baseline on the long-tail set is as high as 19%. In addition, we probe models with different template formats and show that template bias decreases in the long-tail distribution.
Table 1: Examples of inferential knowledge in each domain of LINT, in premise,conclusion format.
Locational
Premise: Person X is located at Long Beach. Conclusion: Person X can plant Palms.
Capability and Advice
Premise: Person X has Asthma. Conclusion: Person X should take Terbutaline.
Temporal
Premise: Plant X vanished in Safavid. Conclusion: Plant X cannot be used to conceal MAS 36.
Natural Properties
Premise: Drawer X has trouble containing Scissors. Conclusion: Router cannot be placed in Drawer X.
LINK is our first step towards the generation of long-tail data. The biggest challenge of generating long-tail data for LLMs is juggling enforcing factual correctness with remaining in lower data likelihood, and our logic-induced knowledge beam search framework provides a satisfying solution to this problem. Our work serves as a starting point of the series of research on systematically generating data in the long-tail distribution, and our dataset is a useful resource for evaluating LLMs in the long-tail distribution.
Figure 2: Overview of knowledge beam search (§ 2.3), consisting of a knowledge model, a critic model and a reranker model. We populate the symbolic rule with values and convert it into a natural language statement. This example demonstrates searching B conditioned on the values of A and Z from previous steps. Note that we only verbalize the predicates containing Person X, because all other predicates contain knowledge that the model should have.
2 Logic-Induced-Knowledge-Search (LINK[Uncaptioned image])
2.1 Long-tail Knowledge Statements
Admittedly, no single language model can resemble the true natural language distribution. However, in Appendix D.2, we show that the long-tail distribution of different language models overlap, and that this evidence supports our assumption that a universal natural language distribution exists, and that one language model can approximate the long-tail distribution of other language models.
Therefore, we approximate the written natural language distribution using the existing largest model that outputs log probability, InstructGPT (text-davinci-003) [^4] Since InstructGPT was deprecated on 2024/01/04, we used llama2-70b for distribution approximation in later studies. See Appendix D.2 for their comparison.. Although it is hard to define the exact boundary between the long-tail and the head distribution, we can define them comparatively: with similar length and format, statements in the long-tail distribution have lower likelihood than those in the head distribution. When we generate long-tail statements, we aim to generate those landing in the lower likelihood distribution, compared to other statements of similar length and format.
Inferential knowledge statements are commonly represented under a premise, conclusion structure (Sap et al., 2019). The premise introduces factual knowledge about a person or an object, and the conclusion is some information about the person or the object deducible from the premise.
Intuitively, for statements with similar length and format, the likelihood of content words in the statements heavily impacts the likelihood of the statements. Therefore, we can create long-tail knowledge statements from a symbolic rule and then fill in content words that are less likely to appear in the training corpus. Creating knowledge statements from symbolic rules has two benefits: (1) Symbolic rules are designed to be correct, so we alleviate the pressure of ensuring the deductive plausibility of the statement throughout the entire generation process. (2) The generation process can be broken down into multiple steps, each of which is conditioned on only one variable. The knowledge generation task will be much easier for the model, and it is easier to maneuver the distribution of individual values than the entire sentence.
2.2 Curating Symbolic Rules
A symbolic rule consists of a premise and a conclusion. The conclusion is a single predicate, while the premise contains a set of predicates connected by &\&& operators. Each predicate is a triple of a verb phrase, a subject and an object, and each variable in the symbolic rule has a designated data type. See example below:
allergic_to(Person X, Allergen A) & one_type_of(Ingredient Z, Allergen A)
& ingredient_in(Ingredient Z, Dish B) $\to$ cannot_eat(Person X, Dish B)
A symbolic rule must follow some constraints to turn into a knowledge statement:
The premise and conclusion need to have the same subject.
Different subjects in the premise and conclusion make the rule more fuzzy and less verifiable.
The symbolic rule should be linearly chained.
Starting from the subject in the premise, we should be able to find a chain of variables that connects the subject to the object in the conclusion (or vice versa), and each variable should appear once and only once. This is to simplify the search process.
The premise and conclusion should not be paraphrases.
For example, allergic_to(Person X, Allergen A) $\to$ reacts_badly_to(Person X, Allergen A) is not a valid symbolic rule. This is to ensure that the symbolic rule contains some reasoning.
The symbolic rule should contain at least 3 variables.
This is to ensure some degree of complexity in the symbolic rule.
The symbolic rule should not contain predicates out of scope of LLMs’ knowledge.
has_height(Tree X, Height Y) is not a valid predicate, because it is unlikely that LLMs have knowledge about the exact height of one tree. This is to avoid hallucination.
We curated symbolic rules that span across four domains: temporal, locational, capability and advice, and natural properties, totaling 149 person-related rules and 268 object-related rules. Details of our symbolic rules are in Appendix A.
2.3 Knowledge Beam Search
Defining search order.
Since all variables are linearly chained, we can search them one by one without repetition. We always start with the subject of the sentence – the person or the object, represented as Datatype X. In the rule in § 2.2, for example, we start with Person X in the premise and find a chain of variables that connects it to the object in the conclusion: X, A, Z, B.
For some rules that call for factual knowledge with only one correct answer, such as age, height, year, etc., we empirically find that it increases the knowledge quality to start from the subject in the conclusion and end with the object in the premise.
Constructing Prompt.
For each variable, we construct a prompt using all predicates that contain that variable and other previously searched variables. For example, to search variable B in the rule in § 2.2, we include predicate ingredient_in(Ingredient Z, Dish B). We assume Z=butter and construct the prompt as follows:
Give me 50 values of B to fill in the sentence “ingredient_in(butter, B)” in the format “1. value.”, where B is a Dish.
Prompting InstructGPT for knowledge.
For each partially searched beam, we obtain 200 values of the current variable from the knowledge model. We call OpenAI API 4 times, generating 50 values each time (temperature=0.7  [^5] We keep top_p=1 for maximum diversity, and top_k is hidden in the OpenAI API. Ablation on temperature is in Appendix C.1.). After each call, we verify the values using a critic model (see paragraph below). To prevent duplicates, we explicitly instruct the model not to generate verified correct values and modify the token probability of incorrect values. We implement an early stop mechanism: if for two consecutive calls we do not get any correct values, we terminate the search for the beam.
Verifying values with a critic.
We use huggingface default implementation of Flan-T5-XXL (Chung et al., 2022), an instruction-tuned model that can be used zero-shot, as the critic that checks data type conformity and factual correctness of the values. We ask the model to output yes/no on the correctness of a given statement. For data type conformity, the statement is “{value} is a {data type}.” For factual correctness, we convert the symbolic predicate into a natural language statement. We obtain the yes token probability and dynamically adjust the threshold for accepting values for different predicates. For details of the critic model, see Appendix B.
Pushing values to long-tail distribution.
At each search step, we convert symbolic predicates into natural language statements (ingredient_in(butter, saag chicken) normal-$\to$ Butter is an ingredient in saag chicken) and concatenate them with “and”. We obtain the sentence likelihood using huggingface default implementation of llama-7B (Touvron et al., 2023a)666llama2 was not released at the time of experiments. and rank the sentences from the lowest likelihood to the highest likelihood. We take top the 75% values unless there are more than 200 values, in which case we take the top 200 values. Then we move on to the next variable. When searching for values in the head distribution, we rank the sentences from the highest to the lowest likelihood.
From 149 person-related rules and 268 object-related rules across four domains, we curate our dataset Logic-Induced-Long-Tail(LINT) that consists of 54K long-tail knowledge statements. In addition, we also release 54K head distribution statements that are also searched with the LINK framework. Domain-wise statistics of LINT can be found in Appendix F.
3 Can LLM generate in the Long-tail Distribution?
3.1 Generating Long-tail Statements from Symbolic Rules
Prompt
In the following sentence, A is a Food allergen, B is a Name of a dish or food, Z is a Ingredient. Find values of A, B, Z to fill in the blank in the sentence ‘If Person X is allergic to [A] and [ Z] is a ingredient in [ B] and [Z] is one type of [ A], then Person X is not able to eat [ B].’ and make it a grammatical and correct sentence. Give me 50 values in the format ‘1. A=, B=, Z=’.
Table 2: An illustration of prompts for zero-shot LLMs, containing a symbolic rule in natural language and its variables with data type specified.
We generate long-tail statements from a subset of 200 symbolic rules from LINT using ChatGPT and GPT4. We prompt the LLMs with a natural language version of the symbolic rule with data types of the variables specified, and ask them to populate the rule with all variables simultaneously. Table 2 shows the corresponding prompt of the symbolic rule example in § 2.2. To ensure long-tailness, we add “Use less frequent terms of A and B and C” in the prompt.  [^7] Other prompts such as “Use terms that are less common” and “Use terms that have lower probability in language model distribution” have similar effect. For each rule, we obtain 200 statements in each of head and long-tail distribution, from ChatGPT and GPT4 respectively.
3.2 LLMs Cannot Reach the Long-tail Distribution with Only Instructions
We calculate the log likelihood over InstructGPT of all statements generated by LINK, ChatGPT and GPT4. We calculate $\delta=mean(D(H))-mean(D(L))$ for each set of statements generated from each symbolic rule, where $D(\cdot)$ means the log likelihood distribution of InstructGPT, $H$ is the set of statements intended to be the head distribution, and $L$ is the set of the statements intended to be in the long-tail distribution.
Figure 3: LINK generations have higher $\delta$ values for most rules, while $\delta$ values of ChatGPT and GPT4 mostly locate around 0.
Figure 4: Only LINK generations fall in the correct distributions on the log likelihood scale of InstructGPT.
Figure 3 illustrates how ChatGPT and GPT4 are not able to generate long-tail statements merely from prompting. Each grid on the x-axis represents a unique symbolic rule, and each grid on the y-axis represents $\delta\in[-1.5,2]$. A $\delta$ close to 0 means that the intended head and long-tail distribution are inseparable, being larger than 0.3 empirically means a decent separation, and being negative means the intended head and long-tail distributions are reverted on the log likelihood scale of InstructGPT.
Averaged across 200 sampled rules, LINT has a positive $\delta$ of 0.48, while ChatGPT and GPT4 each has a delta of -0.14 and -0.02. The $\delta$ values for LINT (blue line) float above 0 for most of the rules, some even being above 0.5. On the other hand, $\delta$ values for ChatGPT (red line) and GPT4 (green line) mostly locate around 0, with many being negative. Low and negative $\delta$ values indicate that LLMs’ generations are in fact not in the long-tail distribution, and sometimes even in the head distribution.
To better illustrate the distribution of statements generated by ChatGPT and GPT4, we plot the log likelihood of the generated statements for one symbolic rule from the three methods Figure 4. To eliminate noise from incorrect statements on the distribution, we only plot the statements that are marked as correct in human evaluation (explained in § 3.3). The likelihood distribution of more rules can be found in Figure 9.
The long-tail statements from LINK clearly fall in a much lower probability distribution than GPT4’s generations. Moreover, GPT4’s generation in the “long-tail distribution” in fact falls in the same distribution as its head distribution.
Ablation on the reranker model.
Rerankers are essential for pushing the generations into the long-tail distribution. Figure 6 presents the distribution comparison of generated statements by LINK and the variant without the reranker. Without the reranker, the generated statements for both head distribution and long-tail distribution are pulled towards the center of the distribution, making them completely inseparable.
Post-hoc reranking over the LLM generated knowledge does not have the same effect as step-wise reranking in LINK. Figure 7 illustrates the distribution of generated statements by LINK, compared to prompt-based GPT4 and prompt-based GPT4 reranked by InstructGPT. Post-hoc reranking barely changed the distribution of generations, even with the same model as the evaluation.
Even though log likelihood is both used by reranker and the evaluation, they are taken from different models and using different inputs. Figure 10 shows that although statements generally tend to fall in the same distribution of llama-7B and InstructGPT, not all rules have cleanly separable generations. In addition, the statements we use for ranking the knowledge beams are shorter than the final statement, as they only consist of partial predicates. Despite these differences, step-wise reranking using a smaller model achieves the separation that post-hoc filtering with the evaluation model cannot achieve.
Our findings highlight the importance of performing variable-level reranking in LINK.
3.3 LLMs Struggle with Data Correctness with Only Instructions
Table 3: LINK has both the highest factual and data type accuracy in human evaluation.
Accuracy ChatGPT GPT4 LINK
Data Type 85.40 91.80 94.23
Factuality 67.50 84.82 88.71
Overall 56.44 78.23 83.95
In addition to distribution correctness, we also evaluate data type conformity and factual correctness of LLMs’ long-tail knowledge generations using crowdworkers from Amazon Mechanic Turk [^8] https://www.mturk.com (AMT). For data type conformity, we ask an AMT worker Is {variable} a {data type}? for each variable in the symbolic rule. For factual correctness, we ask an AMT worker Does the premise entail the conclusion? We sample 4,000 statements from LINT for human evaluation, of which 2,025 are from head distribution and 1,975 are from long-tail distribution. We take 3 annotations for each statement and take the majority vote. Annotator agreement can be found in Appendix G.3. The AMT template can be found in Appendix G.2.
Table 3 shows that prompt-based ChatGPT and GPT4 underperform LINK in both data type conformity and factual correctness. Without LINK, both models struggle more with factual correctness, a foreseeable behavior in the low likelihood realm. For domain wise performance see Table 14. For examples of failure cases, see Table 17.
Ablation on the critic model.
Critic models are essential for guaranteeing the generation quality, especially in the long-tail distribution. Table 8 in Appendix C.2 shows that removing the critic and removing both reranker and critic leads to significant drops of data type conformity and factual correctness of generations in the long-tail distribution. Note that LINK w/o reranker + critic has higher generation quality than LINK w/o critic because the former framework cannot generate long-tail statements without the reranker, so the generations are in the head distribution. This observation further suggests that critic models are essential for generation qualities in the long-tail distribution.
3.4 LINK Facilitates Long-tail Generation Regardless of the Knowledge Model
Our analysis above has shown that by simply prompting ChatGPT and GPT4, arguably the state-of-the-art LLMs on language tasks, we cannot effectively generate knowledge statements that are both high quality and in the long-tail distribution. By adding LINK, we can improve both distribution correctness and generation quality, as shown in Figure 8 and Table 10.
Interestingly, we find that the generations from LINK + GPT4 do not show a big improvement over LINK + InstructGPT, both on distribution correctness and generation quality. This suggests that the improvement a stronger knowledge model brings to LINK is marginal compared to that of the reranker and critic model. This finding highlights the effectiveness of LINK in facilitating long-tail generation regardless of the knowledge model.
Domain Distribution Template InstructGPT ChatGPT GPT4 Human Baseline
Zero-Shot COT Zero-Shot COT Zero-Shot COT
Natural Properties Head Positive Template 3.47 18.52 33.54 35.95 31.33 50.13 -
Negative Template 93.85 70.75 72.68 60.96 89.05 81.58 -
All 52.13 46.64 54.62 49.41 62.41 67.07 82.31
Longtail Positive Template 3.71 16.97 33.39 34.09 24.62 44.53 -
Negative Template 93.84 70.50 71.21 60.13 87.36 79.57 -
All 52.24 45.79 53.75 48.11 58.40 63.40 82.45
Temporal Head Positive Template 12.12 47.96 45.28 59.83 79.43 87.03 -
Negative Template 88.23 79.33 83.57 80.35 86.20 79.81 -
All 53.10 64.85 65.90 70.88 83.07 83.14 84.69
Longtail Positive Template 16.56 47.01 44.29 58.72 77.41 81.99 -
Negative Template 83.89 76.02 79.90 76.92 83.08 76.39 -
All 52.81 62.63 63.46 68.52 80.46 78.97 83.20
Capability and Advice Head Positive Template 22.61 52.85 61.84 70.38 74.64 78.46 -
Negative Template 84.80 71.65 77.66 74.91 86.93 84.63 -
All 56.10 62.97 70.36 72.82 81.26 81.78 83.83
Longtail Positive Template 27.66 55.38 67.01 68.77 72.87 80.89 -
Negative Template 78.22 69.71 74.19 74.24 82.87 79.60 -
All 54.88 63.09 70.88 71.71 78.25 80.19 85.13
Locational Head Positive Template 22.86 52.86 51.67 63.10 52.14 59.29 -
Negative Template 87.55 65.92 68.98 61.84 75.71 77.14 -
All 57.69 59.89 60.99 62.42 64.84 68.90 75.71
Longtail Positive Template 41.39 57.30 51.87 57.68 59.74 63.30 -
Negative Template 76.24 62.12 63.40 58.11 73.03 70.30 -
All 60.16 59.90 58.08 57.91 66.90 67.07 67.42
Table 4: Performance on the entailment classification task of three LLMs decreases on the long-tail distribution compared to the head distribution, while human performance does not. LLMs also have a template bias towards negative templates – when the target token is “no”, “false” and “wrong”.
4 Can LLM Reason Effectively in the Long-tail Distribution?
We used all human evaluated knowledge statements that are labeled correct or factually incorrect in LINT, with 1,925 statements in head distribution and 1,856 statements in long-tail distribution, and created a simple entailment classification task to motivate usage of our LINT dataset.
We evaluate three LLMs: InstructGPT, ChatGPT and GPT4, using both zero-shot and Chain-of-Thought (COT) (Wei et al., 2022) prompting.
We use data labeled as both factually correct and conforming data type as positive samples and use data labeled as factually incorrect during human evaluation as negative samples. For all positive samples, entailment forms between the premise and conclusion, while for negative samples, non-entailment forms. Given that LINT has an 84% accuracy, we have more positive samples than negative samples.
In order to examine whether data distribution affects reasoning behavior in LLMs, we convert each statement into 13 probing questions, where each probing template corresponds to either a positive label (“Yes”, “True”, or “Right”) or a negative label (“No”, “False”, or “Wrong”). The probing templates are summarized in Appendix E.1.
We obtain human performance on the same set of statements. We recruit 17 AMT workers who do not participate in the evaluation task (and thus have not seen the task data). The workers see the knowledge statements in premise, conclusion format and are asked to select “yes/no” to whether the premise entails the conclusion. The workers are asked to use search engines to verify their answers. See AMT templates in Appendix G.2.
For each domain, we report aggregated and template-wise performance of InstructGPT, ChatGPT and GPT4 as well as human baseline performance in Table 4. Random guessing performance is 54% instead of 50% because of the 13 probing templates.
Performance drops in the long-tail distribution.
The performance drop of GPT4 is as high as 4% in zero-shot setting and as high as 5% in COT setting. Moreover, the long-tail performance of models have a marked gap compared to human baselines, being as high as 19% in the “Natural Properties” domain.
The positive template performance experience a “long-tail rise” for all models and most domains.
For all models, the performance on positive templates is significantly lower than on negative templates. Besides “Natural Properties” domain, we witness an increase in performance in the long-tail distribution on the positive templates, contrary to our observation on the negative templates. We examine the rationale the models generate during COT in Table 12 and find that the models tend to avoid drawing a “definite conclusion”. We hypothesize that this phenomenon may be a result of model adjustments made with Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022), and these model adjustments may have a weaker effect on the long-tail distribution due to the knowledge being less frequent during training. This hypothesis can be further supported by the significant increase in positive template performance after COT prompting, where COT weakens the model’s original distribution over the answer tokens in the in-context examples. We leave the exact explanation to future work.
Human Performance does not drop for long-tail distribution.
Performance drop in the long-tail distribution does not occur to humans for 3 out of 4 domains. It is expected because humans can verify their knowledge using search engines, so infrequent knowledge does not challenge humans as much as models. The exception with the locational domain may be due to some relations being less available online(eg. banned_in(Food, Country)).
5 Related Work
The community has begun to realize the importance of testing language models’ abilities in the long-tail distribution (Godbole and Jia, 2022). Recent works reveal that the performance of LLMs is affected by input data probability. (McCoy et al., 2023; Razeghi et al., 2022), and more works have focused on generating less common data for probing LLMs. RICA (Zhou et al., 2020) proposes to include novel entities in self-contained commonsense statements to evaluate robust inference capabilities. UnCommonSense (Arnaout et al., 2022) proposes to evaluate models on informative negative knowledge about everyday concepts in addition to positively expressed commonsense knowledge. Razeghi et al. (2022) observe a correlation between the model performance on math problems and the frequency of numeric and temporal terms from those instances in the pretraining data.
In addition to probing models’ reasoning ability on less common data, recent works also propose to test language models’ ability to generate less common dataChen et al. (2023) propose a negative knowledge generation task where models generate uncommon knowledge with negation conditioned on constrained keywords. Tang et al. (2023) introduce the “less likely brainstorming” task that asks a model to generate outputs that humans think are relevant but less likely to happen.
Generating uncommon data is not only challenging for language models, but also challenging for humans because of our cognitive biasTversky and Kahneman (1974) observe that humans are prone to more systematic errors when facing uncertain events, and Tversky and Kahneman (1973) reveal that humans tend to evaluate the frequency of classes or the probability of events by availability, i.e., by the ease with which relevant instances come to mind. These traits make it difficult for humans to come up with novel associations (Kray et al., 2006), a crucial ability to create data in the long-tail distribution.
6 Conclusion and Future Work
This paper is situated in the problem of generating challenging evaluation data for LLMs in the long-tail distribution. We proposed a long-tail data generation framework, LINK, which grounds the statements in symbolic rule templates. We showed that zero-shot long-tail knowledge generation is still hard for the SOTA LLMs, and our approach LINK is effective for reaching the long-tail distribution while preserving factual correctness. We also showed that LLM’s reasoning capabilities are especially affected in the long-tail regime.
Using LINK, we created a dataset, LINT, which contains 417 symbolic rules and 54K long-tail knowledge statements. LINT is a useful resource for curating challenging evaluation data for LLMs. Using a simple entailment classification task, we showed that the long-tail data reveals more flaws of LLM’s reasoning capabilities.
LINK can facilitate future work on data augmentation. One can take natural language statements from existing benchmarks (or new benchmark proposals) and turn them into symbolic rules, from which we can generate long-tail data using LINK. Given such long-tail data are less frequently seen during training, it is also useful for mitigate benchmark leakage issues during evaluation.
Checklist
Limitation
Limitation on knowledge statement format.
Long-tail knowledge statements may come in multiple shapes and forms. Our work focuses only on premise, conclusion format, as the first step towards the generation of knowledge statements. The symbolic rules do not have high complexity, due to the limited number of variables and predicates, and being under the constraint for the symbolic rules to be linearly chained. Therefore, the effectiveness of our framework on generating more complex knowledge statements has not been tested.
Limitation on testing with open-source models.
Our work did not include open-source models in evaluations of long-tail statement generation and entailment classification task. While ChatGPT and GPT4 are arguably the strongest models, open-source models may exhibit new behaviors in the long-tail realm that are worth exploring.
Limitation on ablating with different critic and reranker model settings.
While we performed extensive ablation studies on the critic and reranker models and established their importance in the LINK framework, we did not explore a diverse set of model options as well as hyperparameter settings. Using other models may or may not affect the performance of LINK.
Limitation on sample size.
Due to constraint from human annotation resources, we were only able to evaluate models on 200 rules uniformly sampled from the LINT. Although the general trend should remain the same, model performance evaluated on all rules may result in some deltas.
Risk
Generation of harmful values.
LINK might be used on mal-intention-ed rules or searching for toxic and harmful values, where researchers may replace our reranker with another model trained to prefer more harmful values.
Environmental tax.
Another potential risk is increasing environmental burdens because we extensitvely call OpenAI APIs to large language models during search; however, one can replace the large language models with smaller open source models with less environmental tax.
Factual errors in generations.
Because LINK operates in the long-tail realm, its generations are no guaranteed to be correct 100% of the time. If one uses the generations directly without verification, one my introduce false information into their system.
Use and Distribution
All data we collected through LLMs in our work are released publicly for usage and have been duly scrutinized by the authors. Data for all human studies that we conduct are also publicly released with this work, with appropriate annotator anonymizations.
Our framework LINK may only be used for generations that follow the ethics guideline of the community. Using LINK on mal-intention-ed rules or searching for toxic and harmful values is a potential threat, but the authors strongly condemn doing so.
Appendix A Symbolic Rule Creation
Following the criteria mentioned in § 2.2, we curated 417 symbolic rules from the following sources:
  1. 1.
    Expansion from an abstract rule. We prompt GPT4 to obtain concrete data types for the abstract rule to create multiple concrete rules. Table 5 shows an example of one abstract rule, related prompts for expansion and one concrete rule. We obtained 201 rules using this method.
  2. 2.
    Direct generation from concrete data type pairs. We prompt GPT4 to generate multiple predicates defining the relationship of the data types as the conclusion, and then generate a premise. We obtained 216 rules using this method.
Table 5: An example meta-rule and its corresponding prompts for expansion.
Meta-Rule
is_of_age(Person A, Age X) & requires_a_ minimal_age_of(Object B, Age Y) & is_smaller_than(Age X, Age Y) $\to$ cannot_operate(Person A, Object B)
Prompt for Data Type Expansion
In rule “requires_a_minimal_age_of(Object B, Age Y) & cannot_operate(Person A, Object B)", B is a variable representing an object. List 10 subcategories of object that B could be that also make the rule true.
Expanded Data Types
Vehicle, Machinery, Alcohol, Firearm, Tattoo Equipment, Tobacco Product
Prompt for Verb Optimization
cannot_operate(Person A, Object B) is equal to [mask](Person A, Vehicle B). Write the best predicate that could fit in [mask] token.
Expanded Rule
is_of_age(Person A, Age X) & requires_a_ minimal_operating_age_of(Object B, Age Y) & is_smaller_than(Age X, Age Y) $\to$ cannot_drive(Person A, Vehicle B)
Appendix B Critic Model
We find that while the critic model usually verifies data type conformity with high accuracy, it often creates false negatives when verifying factual correctness. Moreover, even within false negatives that result from the same predicate, the correct values get higher yes token probabilities than the incorrect values. We hypothesize that while the critic model is less confident about certain knowledge because it is trained on a smaller portion of the knowledge than text-davinvi-003, it can still rank the values inherently. Therefore, we extract the probability of the yes token instead of taking the argmax. We also implement a dynamic critic threshold that adjusts the threshold for accepting values for different predicates. The algorithm is as follows:
  1. 1.
    We start with a threshold of 0.85.
  2. 2.
    If no correct values are found, we decrease the threshold by 0.05.
  3. 3.
    If some correct values are found, we set the threshold for the predicate to the current threshold and do not decrease it in further calls.
  4. 4.
    If the threshold is set but we find some values with a higher yes token probability than the threshold, we increase the threshold by an increment of 0.05 to accommodate the higher probability. Then we retrospectively reject previous accepted values with a lower yes token probability than the new threshold.
  5. 5.
    For data type conformity, we set a minimum threshold of 0.65 because we expect the model to be more confident.
In this way, we can find the maximum available threshold for each beam, which guarantees precision while reducing false negatives.
To verify the effectiveness of our critic model, we use crowd workers from AMT to evaluate the data type conformity and factual correctness of predicates. Specifically, for each symbolic predicate that contains two variables (e.g., exist_during(Location X, Historical Time Period Y)), we will present a statement in natural language (e.g., Saigon existed during The Cold War.) with 3 types of questions: (1) clear reference: Q1 and Q2. (2) factual correctness: Q3. (3) data type conformity: Q4 and Q5.
  • Q1: Does “Value A” in the Statement “Statement” have a clear reference?
  • Q2: Does “Value B” in the Statement “Statement” have a clear reference?
  • Q3: Is the Statement “Statement” factually correct, with very high probability?
  • Q4: Is the Statement “Value A is a Data Type A.” factually correct, with very high probability?
  • Q5: Is the Statement “Value B is a Data Type B.” factually correct, with very high probability?
We sample 3 rules from our data and requested human annotators to rate the data type conformity and factual correctness of statements. Table 6 shows the error rate of each question. Only if all the questions are answered with “Yes” do we consider the statement as correct. The overall correctness of statements in head distribution and long-tail distribution are 0.8567 and 0.8467 respectively, which indicates a high quality of statements accepted by our critic model.
Q1 Q2 Q3 Q4 Q5
Error Rate 0.0004 0 0.0639 0.0011 0
Table 6: The error rate of each question in human verification. Most errors occur on factual correctness.
Appendix C Ablation Studies on LINK
C.1 Hyperparameters on Knowledge Model
When constructing LINT, we used InstructGPT as the knowledge model with temperature=0.7 and top_p=1. Since top_p=1 maximizes sampling diversity and top_k is hidden from the OpenAI API, we conduct ablation studies on whether temperature affects the result of knowledge search, comparing temperature of 0.5 (low diversity), 0.7 (medium diversity) and 1.0 (high diversity), using a few sampled rules. In this ablation study, we use gpt-3.5-turbo-instruct checkpoint as knowledge model and llama-2-70b as the approximation of the language distribution.
Table 7: Different temperatures result in similar data type conformity and factual correctness.
Temperature Data Type Factuality Overall
0.5 89.05 93.21 82.94
0.7 89.00 92.33 82.25
1.0 88.08 91.75 81.00
Table 7 shows similar data type conformity and factual correctness among the three ablated temperature, with temperature=1.0 having the lowest accuracy among the three settings.
Figure 5 shows that all three temperature settings can successfully generate knowledge statements in the long-tail distribution, except for when temperature=0.5 in one of the six sampled rules.
This phenomenon reflects that higher temperature helps generating more diverse values and therefore more likely to generate long-tail values, while risking lowering factual salience.
(a) rule0
(b) rule32
(c) rule46
(d) rule88
(e) rule112
(f) rule122
Figure 5: All three temperature settings of LINK can successfully generate knowledge statements in the long-tail  distribution, except for when temperature=0.5 in one of the six sampled rules.
C.2 Effect of Critic on LINK
To investigate the effectiveness of the critic, we provide an ablation study on a few sampled rules by removing the critic in LINK. Table 8 shows the generation quality of LINK and several variants in long-tail distribution. Without the critic, the generation quality decreases significantly. However, the performance drop is less significant in the head distribution Table 9. Besides, if we replace the reranker with a random sampling method, the generated statements cannot lie in the long-tail distribution (which will be further explained in § C.3) and have higher correctness without the critic. It indicates that it is harder for models to generate correct statements from the long-tail distribution than the head distribution without LINK.
Table 8: Ablation study on the critic model in the long-tail distribution. Removing the critic from LINK will significantly decrease the generation quality. Using a critic is necessary to guarantee the correctness of generated statements, especially in the long-tail distribution.
Data Type Factuality Overall
LINK 93.42 97.50 91.33
  w/o critic 52.58 52.08 33.00
  w/o critic+reranker 75.42 73.92 58.25
LINK with GPT4 92.75 96.17 89.25
  w/o critic 63.00 58.17 40.00
  w/o critic+reranker 88.33 83.50 74.50
Table 9: Ablation study on the critic model in the head distribution. Removing the critic decreases the data quality, but not as much as in the long-tail distribution. LINK w/o critic+reranker has the same performance between head and long-tail distribution, demonstrating that without a reranker all generations are in the same distribution.
Data Type Factuality Overall
LINK 95.17 97.75 93.00
  w/o critic 80.33 73.50 59.75
  w/o critic+reranker 76.66 74.83 59.33
LINK with GPT4 92.17 98.08 90.83
  w/o critic 91.00 80.42 72.08
  w/o critic+reranker 88.17 85.33 75.58
C.3 Effect of Reranker on LINK
To investigate the effectiveness of the reranker, we provide an ablation study on a few sampled rules by replacing the reranker step with a random sampling method. Figure 6 presents the distribution comparison of generated statements by LINK and the variant without the reranker. Without the reranker, the generated statements for both head distribution and long-tail distribution are pulled towards the center of the distribution, making them completely inseparable.
(a) rule0
(b) rule32
(c) rule46
(d) rule88
(e) rule112
(f) rule122
Figure 6: When we remove reranker from LINK, the distribution of the resulting head and long-tail statements are pulled towards the center. Using reranker is essential for separating the head and long-tail distribution.
C.4 Ineffectiveness of Post-hoc Reranking for LLM generated knowledge.
To further highlight the importance of performing step-wise reranking in LINK, we confirm that applying a post-hoc reranker on the GPT4 generations from instructions does not have the same effect as LINK. We use InstructGPT to rerank the GPT4 generations from the lowest to the highest likelihood and take the top 75% results as the long-tail distribution. For the head distribution, we rerank the generations from the highest to the lowest likelihood and take the top 75% results.
Figure 7: Post-hoc reranking of GPT4 does not help move the distribution towards the long-tail distribution.
We evaluate on the same set of rules as in § 3.2 as an example. Figure 7 illustrates the distribution of generated statements by LINK, prompt-based GPT4 and prompt-based GPT4 with reranker. We observe that using post-hoc reranker still cannot achieve a separation between the generation of the head distribution and the long-tail distribution, even with the same model as the evaluation. It demonstrates that maneuvering the distribution during the searching process is necessary and more effective than post-hoc filtering.
C.5 Applying GPT4 as the knowledge model
Table 10 shows the generation quality of GPT4 using baseline prompting method, LINK and LINK with GPT4 as the knowledge model over 6 sample rules. Using a stronger model as the knowledge model has marginal effect on the quality of generations compared to LINK. Figure 8 shows that whatever the knowledge model is, the distribution of generations by LINK can correctly fall in the long-tail distribution.
Table 10: Using a stronger model as the knowledge model does not improve generation qualities for LINK, but using LINK with a language model has significant improvement over zero-shot performance.
Data Type Factuality Overall
Zero-shot GPT4 85.44 88.42 74.39
LINK 93.42 97.50 91.33
LINK with GPT4 92.75 96.17 89.25
(a) rule0
(b) rule32
(c) rule46
(d) rule88
(e) rule112
(f) rule122
Figure 8: LINK using GPT4 creates statements that fall in a roughly similar long-tail distribution as the original LINK with InstructGPT.
Appendix D Addendum on Distribution
(a) rule0
(b) rule32
(c) rule46
(d) rule88
(e) rule112
(f) rule122
Figure 9: An illustration on the distribution of generated statements by LINK, ChatGPT and GPT4. While LINK’s long-tail generations fall into a lower probability distribution than those of GPT4, GPT4’s “long-tail distribution" overlaps with the head distribution, indicating that these generations are not truly long-tail.
D.1 Additional distribution plots for symbolic rules
As an extension on § 3.2, we show the distribution of statements sampled by LINK, ChatGPT and GPT4 from 6 symbolic rules on InstructGPT in Figure 9.
D.2 Distribution Comparison of Different Models
In this section, we show that the long-tail distribution of different language models overlap, and that this evidence supports our assumption that a universal natural language distribution exists; subsequently, the long-tail distribution of a language model can be used to approximate the long-tail distribution of other language models.
We sample knowledge statements generated by LINK from six rules and calculate their probabilities with llama-7b, llama2-7b, llama-2-70b, and InstructGPT. Figure 10, Figure 11 and Figure 12 respectively show the distribution comparison between InstructGPT and the three open-source models over the sampled statements from each rule.
For every rule, we note that if a set of statements falls into the low-probability distribution of InstructGPT, it also falls into the low-probability distribution of the open-source model. Therefore, the categorization on long-tail distribution by one language model can effectively approximate the categorization on long-tail distribution by other models; hence, we use InstructGPT as the approximation of the written natural language distribution in our distribution evaluation.
(a) rule0
(b) rule32
(c) rule46
(d) rule88
(e) rule112
(f) rule122
Figure 10: An illustration of the distribution comparison between llama-7B and InstructGPT of generated statements by LINK.
(a) rule0
(b) rule32
(c) rule46
(d) rule88
(e) rule112
(f) rule122
Figure 11: An illustration of the distribution comparison between llama2-7B and InstructGPT of generated statements by LINK.
(a) rule0
(b) rule32
(c) rule46
(d) rule88
(e) rule112
(f) rule122
Figure 12: An illustration of the distribution comparison between llama2-70B and InstructGPT of generated statements by LINK.
Appendix E Entailment Classification Probing
E.1 Probing Template
Table 11 shows templates we used for the entailment classification task. As mentioned in § 4, we divide the templates into positive templates and negative templates. Positive templates are those with a positive label (i.e., Yes, Right and True) and negative templates are those with a negative label (i.e., No, Wrong and False).
Most of the templates have definite labels across all rules. However, the label of Template 7 depends on the rules. If the rule has a positive conclusion (e.g., Person X can use ChatGPT), the answer to the question should be positive, i.e., Yes. On the contrary, if the rule has a negative conclusion(e.g., Person X cannot use ChatGPT), the answer to the question should be negative, i.e., No.
Table 11: Templates used for machine entailment classification task.
Template Label
1 Is it true that if premise, conclusion. Yes
2 Yes or no: if premise, conclusion. Yes
3 True or false: if premise, conclusion. True
4 Right or Wrong: if premise, conclusion. Right
5 Premise: premise. Conclusion: conclusion. Does premise entail conclusion? Yes
6 Premise: premise. Conclusion: conclusion_negation. Does premise contradict the conclusion? Yes
7 Answer the question with yes or no: if premise, conclusion_question? Depends
8 Is it true that if premise, conclusion_negation. No
9 Yes or no: if premise, conclusion_negation. No
10 True or false: if premise, conclusion_negation. False
11 Right or Wrong: if premise, conclusion_negation. Wrong
12 Premise: premise. Conclusion: conclusion_negation. Does premise entail conclusion? No
13 Premise: premise. Conclusion: conclusion. Does premise contradict the conclusion? No
E.2 Rationale analysis
As mentioned in § 4, we examine the rationale the model generated during COT in the entailment classification task and found that the models tend to avoid drawing a “definite conclusion”. Table 12 shows an example.
Table 12: A rationale generated by GPT4 with COT shows that even though the model is able to cite the relevant information, it inclines to answer “No" because it cannot “definitively conclude".
Input
Premise: PersonX is of age 17 years 8 months. Conclusion: PersonX cannot consume Bourbon. Does premise entail conclusion?
Label
Yes
Answer
No
Rationale
In many countries, the legal drinking age is 18 or 21. If we do not know the specific country PersonX is in, we cannot definitively conclude that they cannot consume Bourbon.
Appendix F Domain-wise statistics of LINT
Table 13 shows the detailed statistics of LINT.
Table 13: The number of symbolic rules and knowledge statements in different domains in LINT.
Rule Head Long-tail
Temporal 81 15,143 15,317
Capability and Advice 132 14,966 15,010
Natural Properties 139 16,788 16,669
Locational 65 7,323 7,370
Total 417 54,220 54,366
F.1 Domain-wise human evaluation
As mentioned in § 3.3, we uniformly sample 4,000 statements from LINT for human evaluation. Table 14 provides more detailed domain-wise statistics on the data type conformity and factual correctness performance of LINT long-tail knowledge generation. While “Natural Properties” has the highest overall accuracy and factuality, model performance on positive templates in Table 4 is the lowest while model performance on negative templates is the highest in this domain. This suggests that these LLMs might have been most aligned in this domain during pre-training.
Table 14: The factual and data type accuracy of each domain in human evaluation.
Data Type Factuality Overall
Temporal 90.18 85.27 77.38
Capability and Advice 94.97 80.73 75.98
Natural Properties 96.61 96.61 93.81
Locational 98.88 70.79 70.79
F.2 Rule definitions
Table 15 shows the definitions of the 6 sampled rules.
Table 15: Rule definitions of six sampled rules.
Rule0
lived_in(Person P, Geographic Location A) & lived_during(Person P, Historical Time Period D) & existed_during(Geographic Location A, Historical Time Period D) & was_invented_in(Product or Technology C, Year Y) & is_more_than_a_century_earlier_ than(Historical Time Period D, Year Y) $\to$ is_not_able_to_use(Person P, Product or Technology C)
Rule32
has_trouble_lifting(Person X, Name of Appliance B) & is_heavier_than(Object A, Name of Appliance B) $\to$ cannot_lift(Person X, Object A)
Rule46
is_allergic_to(Person A, Substance X) & includes(Name of Cosmetics B, Substance X) $\to$ cannot_use(Person A, Name of Cosmetics B)
Rule88
died_in(Historical Figure A, Historical Time Period X) & was_created_during(Artifact B, Historical Time Period Y) & is_earlier_than( Historical Time Period X, Historical Time Period Y) $\to$ cannot_create(Historical Figure A, Artifact B)
Rule112
has_trouble_containing(Room B, Furniture C) & is_larger_than(Furniture A, Furniture C) $\to$ cannot_fit_in(Furniture A, Room B)
Rule122
has_trouble_containing(Trunk B, Furniture C) & is_larger_than(Furniture A, Furniture C) $\to$ cannot_fit_in(Furniture A, Trunk B)
Appendix G Amazon Mechanic Turk
G.1 Recruiting Workers
We recruit workers from all English-speaking countries (US, UK, New Zealand, Australia, Canada), although AMT workers are mostly US-based. We use a qualification task to recruit AMT workers. In the qualification task, all workers will be presented with three manually selected statements, which are clear and representative. Each statement has five related questions as described in Appendix B. Only workers who answer all the questions correctly will be recruited. In the end, we recruited 38 workers to evaluate the quality of generation and 17 workers as human baselines for the entailment classification task. We paid the workers \$0.47 per annotation for evaluating the quality of generations and \$0.11 per annotation for the entailment classification task, to match \$15 per hour based on their working time.
G.2 Templates
Figure 13 and Figure 14 show the template we use for the evaluation of generation quality and the entailment classification task.
G.3 Agreement statistics
Table 16 shows the agreement of annotations in the evaluation task. The high agreement of the data type conformity and factual correctness for LINT ensures the reliability of our results. The agreement for baselines is lower, which also indicates that the generated statements of baselines are of low quality and confusing for human annotators.
Table 16: Agreement of annotations in the evaluation task.
Accuracy ChatGPT GPT4 LINK
Data Type 79.29 83.16 87.54
Factuality 38.35 58.48 75.10
Overall 65.64 74.93 83.39
G.4 Failure Case Examples
We analyze some failure cases that are labeled as incorrect in the human evaluation. Table 17 presents some examples.
Table 17: Examples that are labeled as incorrect during human evaluation. Note that the reasons are analyzed by the authors instead of annotators.
Rule 172
Locational
Rule: is_located_in(Person A, Location X) & is_forbidden_in(Food Item B, Location X) $\to$ cannot_eat(Person A, Food Item B) Premise: Person X is located in Houston Conclusion: Person X cannot eat Chocolate Is Houston a location? Annotation: Yes Is Chocolate a food item? Annotation: Yes Does the premise entail the conclusion? Annotation: No Reason: It is a factual error. Chocolate is not actually forbidden in Houston, so People in Houston can eat chocolate.
Rule 371
Capability and Advice
Rule: can_treat(Drug B, Name of Disease X) & has(Person A, Name of Disease X) $\to$ should_take(Person A, Drug B) Premise: Person X has Hepatitis Conclusion: Person X should take Sofosbuvir Is Hepatitis a name of disease? Annotation: Yes Is Sofosbuvir a drug? Annotation: Yes Does the premise entail the conclusion? Annotation: No Reason: It is a factual error. There are different types of hepatitis viruses. Sofosbuvir is a medication used primarily for the treatment of hepatitis C. For other types of hepatitis, different medications or treatments may be necessary.
Rule 274
Temporal
Rule: vanished_in(Plant A, Historical Time Period X) & was_invented_in(Weapon B, Historical Time Period Y) & is_earlier_than(Historical Time Period X, Historical Time Period Y) $\to$ cannot_be_used_to_conceal(Plant A, Weapon B) Premise: Plant X vanished in Mongol Conclusion: Plant X cannot be used to conceal M92 Zolja Is Mongol a historical time period? Annotation: No Is M92 Zolja a weapon? Annotation: Yes Does the premise entail the conclusion? Annotation: Yes Reason: It is a data type error. The Mongols are an East Asian ethnic group native to Mongolia, not a time period. The Mongol Empire may refer to a period of the 13th and 14th centuries, but Mongol cannot.
Rule 204
Natural Properties
Rule: has_trouble_containing(Drawer B, Tool C) & is_larger_than (Tool A, Tool C) $\to$ cannot_be_placed_in(Tool A, Drawer B) Premise: Drawer X has trouble containing Scroll saw Conclusion: Car cannot be placed in Drawer X Is Scroll saw a Tool? Annotation: Yes Is Car a Tool? Annotation: No Does the premise entail the conclusion? Annotation: Yes Reason: It is a data type error. Car is a vehicle instead of a tool.
Refer to caption Figure 13: AMT template for the evaluation of generation quality. Refer to caption Figure 14: AMT template for the human baseline of the entailment classification task.

### v3

@@
 Abstract
-State-of-the-art LLMs outperform humans on reasoning tasks such as Natural Language Inference.
-Recent works evaluating LLMs note a marked performance drop on input data from the low-probability distribution, i.e., the long-tail.
-Therefore, we focus on systematically generating statements involving long-tail inferential knowledge for more effective evaluation of LLMs in the reasoning space.
-We first propose a novel framework Logic-Induced-Knowledge-Search (LINK[Uncaptioned image]) [^1] https://github.com/INK-USC/LINK/ that generates factually correct and long-tail knowledge statements grounded on symbolic rule templates; LINK effectively generates data in the long-tail distribution that zero-shot prompted LLMs are unable to reach, and outperforms zero-shot GPT4 on factual correctness by 5%.
-We further use the data generated by LINK to construct a dataset Logic-Induced-Long-Tail (LINT) [^2] https://doi.org/10.5281/zenodo.10691770 that can be used to evaluate downstream models on the long-tail distribution; LINT contains 108K knowledge statements spanning four domains.
-We use LINT to test LLMs on an entailment classification task and find that model performances drop by as high as 5% in the long-tail distribution compared to head distribution.
-Our work shows the utility of evaluating models in the long-tail distribution, and calls for more research on generating evaluation data in the long-tail distribution.
+To effectively use large language models (LLMs) for real-world queries, it is imperative that they generalize to the long-tail distribution, i.e., rare examples where models exhibit low confidence.
+In this work, we take the first step towards evaluating LLMs in the long-tail distribution of inferential knowledge.
+We exemplify long-tail evaluation on the Natural Language Inference task.
+First, we introduce Logic-Induced-Knowledge-Search (LINK[Uncaptioned image]), a systematic long-tail data generation framework, to obtain factually-correct yet long-tail inferential statements.
+LINK uses variable-wise prompting grounded on symbolic rules to seek low-confidence statements while ensuring factual correctness.
+We then use LINK to curate Logic-Induced-Long-Tail (LINT), a large-scale long-tail inferential knowledge dataset that contains 108K statements spanning four domains.
+We evaluate popular LLMs on LINT; we find that state-of-the-art LLMs show significant performance drop (21% relative drop for GPT4) on long-tail data as compared to on head distribution data, and smaller models show even more generalization weakness.
+These results further underscore the necessity of long-tail evaluation in developing generalizable LLMs.
+[^1] https://github.com/INK-USC/LINK
 […]
@@
 In Search of the Long-Tail: Systematic Generation of Long-Tail Inferential Knowledge via Logical Rule Guided Search
-[…]
-Figure 1: An illustration of the distribution of statements generated by different methods, after being filtered on factual correctness by human evaluation.
-For the same symbolic rule, all methods generate the same number of statements, but LINK(Refer to caption) has more correct statements than zero-shot ChatGPT and zero-shot GPT4(Refer to caption).
-LINK long-tail generations clearly lie in a more long-tail part of the probability distribution than the zero-shot GPT4 generations.
+Huihan Li1     Yuting Ning2     Zeyi Liao2     Siyuan Wang3     Xiang Lorraine Li4,5 Ximing Lu4,6     Wenting Zhao7     Faeze Brahman4     Yejin Choi4,6     Xiang Ren1,4 1University of Southern California   2 Ohio State University   3Fudan University 4Allen Institute for Artificial Intelligence   5 University of Pittsburgh 6 University of Washington   7 Cornell University {huihanl,xiangren}@usc.edu
+Figure 1: Overview of our motivation, long-tail data generation framework, and model evaluation.
 1 Introduction
-Large language models (LLMs) have approached broadly human-level performance on an increasing number of benchmarks (OpenAI, 2023; Ouyang et al., 2022).
-The problem of evaluating models that potentially outperform (non-expert) humans on most skills relevant to the task seems more imminent than ever (Bowman et al., 2022).
-Natural Language Inference (Bowman et al., 2015; Zellers et al., 2019), which requires extensive knowledge and complex reasoning about entities and events, is one of such tasks on which LLMs have impressive performance (Achiam et al., 2023; Touvron et al., 2023b; Jiang et al., 2023).
-Recent works on evaluating LLMs on challenge reasoning sets notice a marked decline in performance when facing input data from the “long-tail”, i.e., the low-probability part of the natural language’s distribution [^3] See Section 2.1 for discussion on the relationship between the probability distribution of natural language and LLMs.(McCoy et al., 2023; Razeghi et al., 2022).
-To have a well-rounded evaluation of LLMs within the reasoning space, one needs access to inferential knowledge statements (Figure 1) collected systematically from the long-tail to form challenging evaluation sets (i.e., long-tail inferential knowledge).
-However, so far there has not been any systematic framework that targets generating data in the long-tail of probability distribution of language models.
-Crowdsourcing long-tail data is difficult because of human cognitive bias (Tversky and Kahneman, 1973, 1974), and LLMs’ generation in the long-tail distribution is hindered by their pretraining task of “most likely" next token (McCoy et al., 2023).
-In this paper, we propose a logical rule-guided long-tail knowledge generation framework, Logic-Induced-Knowledge-Search (LINK[Uncaptioned image]), for generating inferential knowledge statements that are both factually correct and in the long-tail  – a task that is fundamentally challenging to SoTA LLMs.
-To achieve the two goals simultaneously, we first ground the generation of inferential knowledge statements with symbolic rules, which consist of variables and predicates that specify the relationship between variables.
-The knowledge statements are ensured to be valid as long as the filled values of the variables satisfy all predicates in the rule.
-Doing so significantly simplifies the task from sentence generation to value search with the given rules.
-Specifically, we propose knowledge beam search (Figure 2) to utilize the modularity of symbolic rules by searching for values of the variables one by one in a designated order – an algorithm inspired by sequential beam search.
-At each step, we prompt an LLM for plausible values that satisfy all predicates in which the variable appears.
-The step-wise search design also allows for granular control of the statement’s correctness and likelihood as it incorporates a critic model that checks for data type conformity and factual correctness and a reranker model that ensures that the values fall into the long-tail of the distribution.
-LINK shows clear advantage in generating long-tail knowledge statements compared to prompting LLMs (Figure 1).
-We find that even with symbolic rules presented in the prompt, e.g., by instructing LLM to fill in values for the rules, the prompt-based methods using ChatGPT (gpt-3.5-turbo) and GPT4 (gpt-4) still struggle to generate knowledge statements that are both correct and in the long-tail distribution.
-This implies that the prompt-based methods may suffer from directly retrieving long-tail knowledge from the natural language distribution.
-In addition, human annotators find that compared to LINK, the statements generated by prompt-based methods are prone to have more factual errors.
-Further analysis shows (1) the key contribution to the long-tail generation quality comes from the LINK pipeline, (2) LINK can be effectively and flexibly applied across different knowledge models, and (3) post-hoc reranking does not reach into the long-tail distribution as LINK.
-Using LINK, we curate a dataset, Logic-Induced-Long-Tail (LINT), which contains 108k knowledge statements spanning across four different domains, presented in a premise, conclusion format (Table 1).
-LINT can be a useful resource for the community to more thoroughly test LLMs’ capabilities in the long-tail distribution.
-We use a subset of LINT to create an evaluation task in the form of entailment classification, where LLMs are asked to judge true or false on a knowledge statement.
-In the long-tail test set, GPT4’s capability in identifying incorrect knowledge drop by as high as 3% with zero-shot prompting, and as high as 4% with Chain-of-Thought prompting (Wei et al., 2022).
-GPT4’s performance gap from human baseline on the long-tail set is as high as 19%.
-In addition, we probe models with different template formats and show that template bias decreases in the long-tail distribution.
-Table 1: Examples of inferential knowledge in each domain of LINT, in premise,conclusion format.
-Locational
-Premise: Person X is located at Long Beach. Conclusion: Person X can plant Palms.
-Capability and Advice
-Premise: Person X has Asthma. Conclusion: Person X should take Terbutaline.
-Temporal
-Premise: Plant X vanished in Safavid. Conclusion: Plant X cannot be used to conceal MAS 36.
-Natural Properties
-Premise: Drawer X has trouble containing Scissors. Conclusion: Router cannot be placed in Drawer X.
-LINK is our first step towards the generation of long-tail data.
-The biggest challenge of generating long-tail data for LLMs is juggling enforcing factual correctness with remaining in lower data likelihood, and our logic-induced knowledge beam search framework provides a satisfying solution to this problem.
-Our work serves as a starting point of the series of research on systematically generating data in the long-tail distribution, and our dataset is a useful resource for evaluating LLMs in the long-tail distribution.
-Figure 2: Overview of knowledge beam search (§ 2.3), consisting of a knowledge model, a critic model and a reranker model.
-We populate the symbolic rule with values and convert it into a natural language statement.
-This example demonstrates searching B conditioned on the values of A and Z from previous steps.
-Note that we only verbalize the predicates containing Person X, because all other predicates contain knowledge that the model should have.
+Generalization, especially to unfamiliar and novel situations, is a cornerstone for the usability of large language models (LLMs) in addressing varied real-world inquiries.
+This imminent demand necessitates evaluation of LLMs in the long-tail  distribution (the space consisting of unfamiliar examples on which the model has low confidence).
+Previous works, mostly focusing on model memorization issue, define long-tail knowledge using the frequency of entities in a knowledge base (Cao et al., 2020), in the pre-training dataset (Kandpal et al., 2023), or in Wikipedia search (Mallen et al., 2022).
+Godbole and Jia (2022) introduces a general definition for long-tail statements, where long-tail examples are assigned lower likelihood by a pre-trained language model.
+We follow this definition which applies to various data format and language task – for any set of statements with similar length and format, those in the long-tail distribution cannot be generated or are generated with low confidence by the models, compared to those in the head distribution.
+Recent works have noticed that LLMs have a marked decline in performance when facing inputs from the long-tail  McCoy et al.
+(2023); Razeghi et al.
+(2022).
+Hallucination, for example, is found to be correlated with data being in the long-tail distribution (Li et al., 2024; Yu et al., 2024).
+LLMs’ ineffective utilization of long-tail knowledge impacts its reasoning capabilities and raises reliability concerns in downstream tasks (Huang et al., 2023).
+Evaluation in the long-tail distribution requires systematic generation of long-tail data.
+However, obtaining examples in the long-tail is non-trivial.
+With state-of-the-art LLMs being trained on vast volume of data on the internet (OpenAI, 2023; Touvron et al., 2023b), it is increasingly difficult to find unseen examples that can effectively test model generalization to its low-confidence end.
+Crowdsourcing long-tail data is also difficult because of human cognitive bias (Tversky and Kahneman, 1973, 1974), and LLMs’ generation in the long-tail distribution is hindered by their pretraining task of “most likely" next token (McCoy et al., 2023).
+While demonstrating long-tail evaluation across all applications and domains is not feasible within the scope of one paper, this work focuses on inferential knowledge statement in the form of Natural Language Inference (NLI) task (Bowman et al., 2015; Zellers et al., 2019): NLI requires extensive knowledge and complex reasoning about entities and events, and is one of such tasks on which LLMs have impressive performance (Achiam et al., 2023; Touvron et al., 2023b; Jiang et al., 2023).
+Following Sap et al.
+(2019), we structure inferential knowledge as if-then relations with variables, written in a premise, conclusion format (Table 1).
+First, we make long-tail inferential knowledge generation possible.
+We propose a novel and lightweight long-tail inferential knowledge generation framework, Logic-Induced-Knowledge-Search (LINK[Uncaptioned image]) (§ 2), a variable-wise prompting framework grounded on symbolic rules.
+This framework enables us to obtain long-tail knowledge statements from existing LLMs.
+Our evaluation shows that by taking simply instructions ChatGPT(gpt-3.5-turbo) and GPT4(gpt-4) can only produce statements in the head distribution that also have lower factual correctness, but using LINK with the LLMs improves on both distribution conformity and factual correctness (§ 3).
+Table 1: Examples of inferential knowledge in each domain of LINT, in a premise (P), conclusion (C) format.
+Locational Head P: Organization X has a branch in Great Lakes Region. C: Organization X has office in North America.
+Long-tail P: Organization X has a branch in Tarapaca Region. C: Organization X has office in South America.
+Outcome and Effect Head P: Person X has Asthma. C: Person X should take Inhaled antiinflammatory drugs.
+Long-tail P: Person X has Hepatitis. C: Person X should take Sofosbuvir.
+Temporal Head P: Plant X vanished in Paleolithic Era. C: Plant X cannot surround Notre Dame de Paris.
+Long-tail P: Plant X vanished in Classical Greece. C: Plant X cannot surround Belém Tower.
+Natural Properties Head P: Bag X has trouble containing Clarinet. C: Upright Piano cannot fit in Bag X.
+Long-tail P: Bag X has trouble containing Pandeiro. C: Dhak cannot fit in Bag X.
+Second, we test LLMs’ long-tail generalization capability on data generated by LINK(§ 4).
+We produce a large-scale dataset, Logic-Induced-Long-Tail (LINT), which contains 108k knowledge statements spanning across 4 different domains(Table 1).
+In the long-tail test set of LINT, GPT4’s capability in identifying incorrect knowledge drop by 21% relative to the test set in the head distribution, and the gap is even larger for other models we tested (ChatGPT, llama2-70b).
+At the same time, human performance significantly outperforms LLMs in both distributions, and stays consistent between head and long-tail test set.
+Our work is the first to propose a systematic framework that generates data in the long-tail distribution.
+Using NLI as an example, we show that generating data in the long-tail distribution is an effective way for curating evaluation examples for LLM generalization.
+Our work serves as a starting point for the series of research on long-tail data discovery and generation, and motivates the community to incorporate long-tail evaluation into model building pipelines.
+Figure 2: Overview of knowledge beam search (§ 2.3).
+We demonstrate searching B conditioned on the values of A and Z from previous steps.
+We only verbalize the predicates containing Person X in the final statement as all other predicates contain knowledge that the model should have.
 2 Logic-Induced-Knowledge-Search (LINK[Uncaptioned image])
-2.1 Long-tail Knowledge Statements
-Admittedly, no single language model can resemble the true natural language distribution.
-However, in Appendix D.2, we show that the long-tail distribution of different language models overlap, and that this evidence supports our assumption that a universal natural language distribution exists, and that one language model can approximate the long-tail distribution of other language models.
-Therefore, we approximate the written natural language distribution using the existing largest model that outputs log probability, InstructGPT (text-davinci-003) [^4] Since InstructGPT was deprecated on 2024/01/04, we used llama2-70b for distribution approximation in later studies.
-See Appendix D.2 for their comparison..
-Although it is hard to define the exact boundary between the long-tail and the head distribution, we can define them comparatively: with similar length and format, statements in the long-tail distribution have lower likelihood than those in the head distribution.
-When we generate long-tail statements, we aim to generate those landing in the lower likelihood distribution, compared to other statements of similar length and format.
-Inferential knowledge statements are commonly represented under a premise, conclusion structure (Sap et al., 2019).
-The premise introduces factual knowledge about a person or an object, and the conclusion is some information about the person or the object deducible from the premise.
-Intuitively, for statements with similar length and format, the likelihood of content words in the statements heavily impacts the likelihood of the statements.
-Therefore, we can create long-tail knowledge statements from a symbolic rule and then fill in content words that are less likely to appear in the training corpus.
-Creating knowledge statements from symbolic rules has two benefits: (1) Symbolic rules are designed to be correct, so we alleviate the pressure of ensuring the deductive plausibility of the statement throughout the entire generation process.
+In this section, we first explain the advantages of generating inferential knowledge grounded on symbolic rules, then illustrate our process of curating symbolic rules, and lastly explain knowledge beam search, our novel variable-wise search pipeline.
+2.1 Advantages of Generating Inferential Knowledge from Symbolic Rules
+Due to the fact that LLMs are pretrained with the task of generating the “most likely" next token, it is fundamentally challenging for them to directly generate long-tail data through prompting that are factually correct and have low likelihood.
+Using symbolic rules to guide the generation of knowledge statements have three benefits: (1) Symbolic rules are designed to be correct, so we alleviate the pressure of ensuring the deductive plausibility of the statement throughout the entire generation process.
 (2) The generation process can be broken down into multiple steps, each of which is conditioned on only one variable.
-The knowledge generation task will be much easier for the model, and it is easier to maneuver the distribution of individual values than the entire sentence.
+Generating for one variable at a time will be much easier for the model, and it is easier to manipulate the distribution of individual values than the entire sentence.
+(3) From one symbolic rule, one can get abundant combinations of variable values as long as they satisfy each predicate in the rule, making the generation process scalable.
 2.2 Curating Symbolic Rules
@@
 Each predicate is a triple of a verb phrase, a subject and an object, and each variable in the symbolic rule has a designated data type.
-See example below:
-allergic_to(Person X, Allergen A) & one_type_of(Ingredient Z, Allergen A)
-& ingredient_in(Ingredient Z, Dish B) $\to$ cannot_eat(Person X, Dish B)
-A symbolic rule must follow some constraints to turn into a knowledge statement:
-The premise and conclusion need to have the same subject.
-Different subjects in the premise and conclusion make the rule more fuzzy and less verifiable.
-The symbolic rule should be linearly chained.
-Starting from the subject in the premise, we should be able to find a chain of variables that connects the subject to the object in the conclusion (or vice versa), and each variable should appear once and only once.
-This is to simplify the search process.
-The premise and conclusion should not be paraphrases.
+While there are infinitely many ways to construct symbols rules, we create ours using the principles of Compatibility and Mutual Exclusivity.
+Compatibility refers to a scenario in which one or more events enables another event to occur.
+We construct the premise and conclusion such that if all predicates in the premise are true, then the predicate in the conclusion can occur.
+Mutual Exclusivity refers to a scenario in which two or more events or conditions cannot occur simultaneously.
+We construct the premise and conclusion such that if all predicates in the premise are true, then the predicate in the conclusion will not occur.
+To construct such conditions, we add constraints such as time, location, or outcomes to variables in the symbolic rules to make their interaction possible/desirable or impossible/undesirable.
+In other words, the conclusion describes an event (certain interaction between the variables) and the premise depicts some combination of conditions.
+For example, a symbolic rule constrained by compatibility looks like:
+exists(Person X, Location  A) & lives(Plant Y, Climate  B)
+& has_climate(Location  A, Climate B)
+$\to$ can_plant(Person X, Plant Y)
+And a symbolic rule constrained by mutual exclusivity looks like:
+exists(Person X, Time Period  A) & exists(Plant Y, Time Period  B) & is_much_later_than(Time Period  A, Time Period B) $\to$ cannot_plant(Person X, Plant Y)
+Here are some additional properties of our symbolic rules:
+No tautologies.
 For example, allergic_to(Person X, Allergen A) $\to$ reacts_badly_to(Person X, Allergen A) is not a valid symbolic rule. This is to ensure that the symbolic rule contains some reasoning.
@@
 has_height(Tree X, Height Y) is not a valid predicate, because it is unlikely that LLMs have knowledge about the exact height of one tree. This is to avoid hallucination.
-We curated symbolic rules that span across four domains: temporal, locational, capability and advice, and natural properties, totaling 149 person-related rules and 268 object-related rules.
-Details of our symbolic rules are in Appendix A.
+We create symbolic rules that span across four domains (of constraint type): temporal, locational, outcome and effect, and natural properties, totaling 149 person-related rules and 268 object-related rules.
+More about symbolic rules in Appendix B.
 2.3 Knowledge Beam Search
@@
 We always start with the subject of the sentence – the person or the object, represented as Datatype X.
-In the rule in § 2.2, for example, we start with Person X in the premise and find a chain of variables that connects it to the object in the conclusion: X, A, Z, B.
+In the rule in Table 2, for example, we start with Person X in the premise and find a chain of variables that connects it to the object in the conclusion: X, A, Z, B.
 For some rules that call for factual knowledge with only one correct answer, such as age, height, year, etc., we empirically find that it increases the knowledge quality to start from the subject in the conclusion and end with the object in the premise.
@@
 For each variable, we construct a prompt using all predicates that contain that variable and other previously searched variables.
-For example, to search variable B in the rule in § 2.2, we include predicate ingredient_in(Ingredient Z, Dish B).
+For example, to search variable B in the rule in Table 2, we include predicate ingredient_in(Ingredient Z, Dish B).
 We assume Z=butter and construct the prompt as follows:
 Give me 50 values of B to fill in the sentence “ingredient_in(butter, B)” in the format “1. value.”, where B is a Dish.
-Prompting InstructGPT for knowledge.
-For each partially searched beam, we obtain 200 values of the current variable from the knowledge model.
-We call OpenAI API 4 times, generating 50 values each time (temperature=0.7  [^5] We keep top_p=1 for maximum diversity, and top_k is hidden in the OpenAI API.
-Ablation on temperature is in Appendix C.1.).
+Prompting for knowledge.
+For each partially searched beam, we obtain 200 values of the current variable from the knowledge model  [^2] We use text-davinci-003 but one can use any model..
+We call OpenAI API 4 times, generating 50 values each time (temperature=0.7  [^3] We keep top_p=1 for maximum diversity, and top_k is unchangable.
+Ablation on temperature in Appendix D.1.).
 After each call, we verify the values using a critic model (see paragraph below).
-To prevent duplicates, we explicitly instruct the model not to generate verified correct values and modify the token probability of incorrect values.
+To prevent duplicates, we explicitly instruct the model not to generate verified correct values and set logit_bias=-100 for incorrect values.
 We implement an early stop mechanism: if for two consecutive calls we do not get any correct values, we terminate the search for the beam.
@@
 We obtain the yes token probability and dynamically adjust the threshold for accepting values for different predicates.
-For details of the critic model, see Appendix B.
-Pushing values to long-tail distribution.
-At each search step, we convert symbolic predicates into natural language statements (ingredient_in(butter, saag chicken) normal-$\to$ Butter is an ingredient in saag chicken) and concatenate them with “and”.
-We obtain the sentence likelihood using huggingface default implementation of llama-7B (Touvron et al., 2023a)666llama2 was not released at the time of experiments.
-and rank the sentences from the lowest likelihood to the highest likelihood.
+More about the critic model in Appendix C.
+Pushing values to long-tail distribution with reranking.
+At each search step, we convert symbolic predicates into natural language statements (ingredient_in(butter, saag chicken) $\to$ Butter is an ingredient in saag chicken) and concatenate them with “and”.
+We obtain the sentence likelihood using huggingface default implementation of llama-7B (Touvron et al., 2023a)444llama2 was not released at the time of experiments.
+and rerank the sentences from the lowest likelihood to the highest likelihood.
 We take top the 75% values unless there are more than 200 values, in which case we take the top 200 values.
 Then we move on to the next variable.
-When searching for values in the head distribution, we rank the sentences from the highest to the lowest likelihood.
+To control data distribution during evaluation, we also generate statements in the head distribution, by ranking the sentences from the highest to the lowest likelihood.
 From 149 person-related rules and 268 object-related rules across four domains, we curate our dataset Logic-Induced-Long-Tail(LINT) that consists of 54K long-tail knowledge statements.
-In addition, we also release 54K head distribution statements that are also searched with the LINK framework.
-Domain-wise statistics of LINT can be found in Appendix F.
-3 Can LLM generate in the Long-tail Distribution?
-3.1 Generating Long-tail Statements from Symbolic Rules
-Prompt
-In the following sentence, A is a Food allergen, B is a Name of a dish or food, Z is a Ingredient.
+We also release 54K head distribution statements that are also searched with the LINK framework.
+Domain-wise statistics of LINT in Appendix H.
+3 Generating Long-tail Inferential Knowledge with LINK
+In this section, we compare LINK’s ability to generate long-tail inferential knowledge with instruction-only LLMs, ChatGPT and GPT4, who do not use knowledge beam search.
+3.1 Instruction-Only Knowledge Generation
+Symbolic Rule    is_allergic_to(Person X, Food allergen A)& is_ ingredient_in(Ingredient Z, Name of a dish or food B) & is_one_type_of(Ingredient Z, Food allergen A) $\to$ is_not_able_to_eat(Person P, Name of a dish or food B)
+Prompt In the following sentence, A is a Food allergen, B is a Name of a dish or food, Z is a Ingredient.
 Find values of A, B, Z to fill in the blank in the sentence ‘If Person X is allergic to [A] and [ Z] is a ingredient in [ B] and [Z] is one type of [ A], then Person X is not able to eat [ B].’ and make it a grammatical and correct sentence.
@@
 Table 2: An illustration of prompts for zero-shot LLMs, containing a symbolic rule in natural language and its variables with data type specified.
-We generate long-tail statements from a subset of 200 symbolic rules from LINT using ChatGPT and GPT4.
-We prompt the LLMs with a natural language version of the symbolic rule with data types of the variables specified, and ask them to populate the rule with all variables simultaneously.
-Table 2 shows the corresponding prompt of the symbolic rule example in § 2.2.
-To ensure long-tailness, we add “Use less frequent terms of A and B and C” in the prompt.
-[^7] Other prompts such as “Use terms that are less common” and “Use terms that have lower probability in language model distribution” have similar effect.
-For each rule, we obtain 200 statements in each of head and long-tail distribution, from ChatGPT and GPT4 respectively.
-3.2 LLMs Cannot Reach the Long-tail Distribution with Only Instructions
-We calculate the log likelihood over InstructGPT of all statements generated by LINK, ChatGPT and GPT4.
-We calculate $\delta=mean(D(H))-mean(D(L))$ for each set of statements generated from each symbolic rule, where $D(\cdot)$ means the log likelihood distribution of InstructGPT, $H$ is the set of statements intended to be the head distribution, and $L$ is the set of the statements intended to be in the long-tail distribution.
+We generate knowledge statements from a subset of 200 symbolic rules from LINT using ChatGPT and GPT4 by only providing it with an instruction.
+We prompt the LLMs with a natural language version of the symbolic rule with data types of the variables specified, and ask them to populate the rule with all variables simultaneously(Table 2).
+Generations from this prompt serve as the head distribution baseline, to which we compare model generations when they are instructed to generate in the long-tail distribution.
+To instruct models to generate long-tail knowledge from symbolic rules, we append “Use less frequent terms of A and B and C” in the prompt.
+[^5] We tried 10 prompts as shown in Appendix F and they have similar effect on model behavior.
+For each rule, we obtain 200 statements using default instruction and 200 statements using long-tail instruction, from ChatGPT and GPT4 respectively. We present our findings below.
+3.2 LINK Generations Consistently Fall in Long-tail Distribution
+Following Godbole and Jia (2022)’s definition of long-tail statements, we use the most capable LLM that was producing log likelihood at the time of experiments (text-davinci-003) to assign likelihood to generated data.
+We compare how different models assign distributions to the same statements, in order to ensure that the distribution stays consistent among all models despite the absolute log likelihood difference (Appendix E.2).
+We calculate the log likelihood over InstructGPT of all statements generated by LINK, compared with instruction-only ChatGPT and GPT4.
+We calculate $\delta=mean(D(H))-mean(D(L))$ for each set of statements generated from each symbolic rule, where $D(\cdot)$ means the log likelihood distribution of the probability model, $H$ is the set of statements as head distribution baseline, and $L$ is the set of the statements intended to be in the long-tail distribution.
+For a model to successfully generate long-tail knowledge statements, the “long-tail” set of sentences should be assigned distinguishably lower probabilities than the sentence in the head distribution.
 Figure 3: LINK generations have higher $\delta$ values for most rules, while $\delta$ values of ChatGPT and GPT4 mostly locate around 0.
 Figure 4: Only LINK generations fall in the correct distributions on the log likelihood scale of InstructGPT.
-Figure 3 illustrates how ChatGPT and GPT4 are not able to generate long-tail statements merely from prompting.
+Figure 3 illustrates that while ChatGPT and GPT4 are not able to generate long-tail statements merely from prompting, LINK is able to generate long-tail statements with much lower likelihood.
 Each grid on the x-axis represents a unique symbolic rule, and each grid on the y-axis represents $\delta\in[-1.5,2]$.
-A $\delta$ close to 0 means that the intended head and long-tail distribution are inseparable, being larger than 0.3 empirically means a decent separation, and being negative means the intended head and long-tail distributions are reverted on the log likelihood scale of InstructGPT.
-Averaged across 200 sampled rules, LINT has a positive $\delta$ of 0.48, while ChatGPT and GPT4 each has a delta of -0.14 and -0.02.
-The $\delta$ values for LINT (blue line) float above 0 for most of the rules, some even being above 0.5.
+A $\delta$ close to 0 means that the intended long-tail distribution generations have the same log likelihoods as the statements in the head distribution, being larger than 0.3 empirically means a decent drop in likelihood, and being negative means the intended long-tail distribution data have even high likelihood than the head distribution data.
+Averaged across 200 sampled rules, LINK has a positive $\delta$ of 0.48, while ChatGPT and GPT4 each has a delta of -0.14 and -0.02.
+The $\delta$ values for LINK (blue line) float above 0 for most of the rules, some even being above 0.5.
 On the other hand, $\delta$ values for ChatGPT (red line) and GPT4 (green line) mostly locate around 0, with many being negative.
-Low and negative $\delta$ values indicate that LLMs’ generations are in fact not in the long-tail distribution, and sometimes even in the head distribution.
-To better illustrate the distribution of statements generated by ChatGPT and GPT4, we plot the log likelihood of the generated statements for one symbolic rule from the three methods Figure 4.
+To better illustrate the distribution of statements generated by LINK, ChatGPT and GPT4, we plot the log likelihood of the generated statements for one symbolic rule from the three methods Figure 4.
 To eliminate noise from incorrect statements on the distribution, we only plot the statements that are marked as correct in human evaluation (explained in § 3.3).
 The likelihood distribution of more rules can be found in Figure 9.
-The long-tail statements from LINK clearly fall in a much lower probability distribution than GPT4’s generations.
-Moreover, GPT4’s generation in the “long-tail distribution” in fact falls in the same distribution as its head distribution.
-Ablation on the reranker model.
-Rerankers are essential for pushing the generations into the long-tail distribution.
-Figure 6 presents the distribution comparison of generated statements by LINK and the variant without the reranker.
-Without the reranker, the generated statements for both head distribution and long-tail distribution are pulled towards the center of the distribution, making them completely inseparable.
-Post-hoc reranking over the LLM generated knowledge does not have the same effect as step-wise reranking in LINK.
-Figure 7 illustrates the distribution of generated statements by LINK, compared to prompt-based GPT4 and prompt-based GPT4 reranked by InstructGPT.
-Post-hoc reranking barely changed the distribution of generations, even with the same model as the evaluation.
-Even though log likelihood is both used by reranker and the evaluation, they are taken from different models and using different inputs.
-Figure 10 shows that although statements generally tend to fall in the same distribution of llama-7B and InstructGPT, not all rules have cleanly separable generations.
-In addition, the statements we use for ranking the knowledge beams are shorter than the final statement, as they only consist of partial predicates.
-Despite these differences, step-wise reranking using a smaller model achieves the separation that post-hoc filtering with the evaluation model cannot achieve.
-Our findings highlight the importance of performing variable-level reranking in LINK.
-3.3 LLMs Struggle with Data Correctness with Only Instructions
+The long-tail statements from LINK clearly fall in a much lower probability distribution than GPT4’s “long-tail” generations.
+Moreover, GPT4’s generation in the “long-tail distribution” in fact falls in the same probability distribution as its head distribution generations.
+3.3 LINK Achieves Higher Data Correctness than Instruction-Only LLMs
 Table 3: LINK has both the highest factual and data type accuracy in human evaluation.
@@
 Overall 56.44 78.23 83.95
-In addition to distribution correctness, we also evaluate data type conformity and factual correctness of LLMs’ long-tail knowledge generations using crowdworkers from Amazon Mechanic Turk [^8] https://www.mturk.com (AMT).
+In addition to distribution correctness, we also evaluate data type conformity and factual correctness of LLMs’ long-tail knowledge generations using crowdworkers from Amazon Mechanic Turk (AMT).
 For data type conformity, we ask an AMT worker Is {variable} a {data type}?
@@
 We take 3 annotations for each statement and take the majority vote.
-Annotator agreement can be found in Appendix G.3.
-The AMT template can be found in Appendix G.2.
-Table 3 shows that prompt-based ChatGPT and GPT4 underperform LINK in both data type conformity and factual correctness.
+Annotator agreement can be found in Appendix I.3.
+The AMT template can be found in Appendix I.2.
+Table 3 shows that instruction-only ChatGPT and GPT4 underperform LINK in both data type conformity and factual correctness.
 Without LINK, both models struggle more with factual correctness, a foreseeable behavior in the low likelihood realm.
@@
 For examples of failure cases, see Table 20.
+3.4 Ablation Studies
+Our results above show that LINK is better than instruct-only generations with ChatGPT and GPT4, in terms of both “long-tailedness” and factual correctness.
+In this section, we perform ablation studies on the role of reranker, critic, and knowledge models in LINK.
+Ablation on reranker.
+Our variable-wise reranker is essential for pushing LINK generations into the long-tail distribution.
+Figure 6 presents the distribution comparison of generated statements by LINK and the variant without the reranker.
+Without the reranker, the statements for the long-tail distribution are pulled towards the head distribution, making the two completely inseparable.
+Post-hoc reranking over LLM generations does not have the same effect as variable-wise reranking in LINK.
+Figure 7 illustrates the distribution of generated statements by LINK, compared to instruction-only GPT4 and instruction-only GPT4 reranked by InstructGPT.
+Post-hoc reranking barely changes the distribution of generations, even when using the same model as the evaluation.
+Even though log likelihood is both used by reranker and evaluation, they are taken from different models and using different inputs.
+The statements we use for ranking the knowledge beams are shorter than the final statement, as they only consist of partial predicates.
+Despite these differences, variable-wise reranking that uses a smaller model achieves the separation that post-hoc filtering with the evaluation model cannot achieve.
+Our findings highlight the importance of performing variable-wise reranking in LINK.
 Ablation on the critic model.
 Critic models are essential for guaranteeing the generation quality, especially in the long-tail distribution.
-Table 8 in Appendix C.2 shows that removing the critic and removing both reranker and critic leads to significant drops of data type conformity and factual correctness of generations in the long-tail distribution.
-Note that LINK w/o reranker + critic has higher generation quality than LINK w/o critic because the former framework cannot generate long-tail statements without the reranker, so the generations are in the head distribution.
+Table 8 in Appendix D.2 shows that removing the critic and removing both reranker and critic leads to significant drops of data type conformity and factual correctness of generations in the long-tail distribution.
+Note that LINK w/o reranker + critic has higher generation quality than LINK w/o critic.
+This is because without a reranker the model is only able to generate statements in the head distribution, making it easier to be factually correct than in the long-tail distribution.
 This observation further suggests that critic models are essential for generation qualities in the long-tail distribution.
-3.4 LINK Facilitates Long-tail Generation Regardless of the Knowledge Model
-Our analysis above has shown that by simply prompting ChatGPT and GPT4, arguably the state-of-the-art LLMs on language tasks, we cannot effectively generate knowledge statements that are both high quality and in the long-tail distribution.
+Ablation on Knowledge Model.
+Our analysis above has shown that by simply providing an instruction to ChatGPT and GPT4, we cannot effectively generate knowledge statements that are both high quality and in the long-tail distribution.
 By adding LINK, we can improve both distribution correctness and generation quality, as shown in Figure 8 and Table 10.
@@
 This finding highlights the effectiveness of LINK in facilitating long-tail generation regardless of the knowledge model.
-Domain Distribution Template InstructGPT ChatGPT GPT4 Human Baseline
-Zero-Shot COT Zero-Shot COT Zero-Shot COT
-Natural Properties Head Positive Template 3.47 18.52 33.54 35.95 31.33 50.13 -
-Negative Template 93.85 70.75 72.68 60.96 89.05 81.58 -
-All 52.13 46.64 54.62 49.41 62.41 67.07 82.31
-Longtail Positive Template 3.71 16.97 33.39 34.09 24.62 44.53 -
-Negative Template 93.84 70.50 71.21 60.13 87.36 79.57 -
-All 52.24 45.79 53.75 48.11 58.40 63.40 82.45
-Temporal Head Positive Template 12.12 47.96 45.28 59.83 79.43 87.03 -
-Negative Template 88.23 79.33 83.57 80.35 86.20 79.81 -
-All 53.10 64.85 65.90 70.88 83.07 83.14 84.69
-Longtail Positive Template 16.56 47.01 44.29 58.72 77.41 81.99 -
-Negative Template 83.89 76.02 79.90 76.92 83.08 76.39 -
-All 52.81 62.63 63.46 68.52 80.46 78.97 83.20
-Capability and Advice Head Positive Template 22.61 52.85 61.84 70.38 74.64 78.46 -
-Negative Template 84.80 71.65 77.66 74.91 86.93 84.63 -
-All 56.10 62.97 70.36 72.82 81.26 81.78 83.83
-Longtail Positive Template 27.66 55.38 67.01 68.77 72.87 80.89 -
-Negative Template 78.22 69.71 74.19 74.24 82.87 79.60 -
-All 54.88 63.09 70.88 71.71 78.25 80.19 85.13
-Locational Head Positive Template 22.86 52.86 51.67 63.10 52.14 59.29 -
-Negative Template 87.55 65.92 68.98 61.84 75.71 77.14 -
-All 57.69 59.89 60.99 62.42 64.84 68.90 75.71
-Longtail Positive Template 41.39 57.30 51.87 57.68 59.74 63.30 -
-Negative Template 76.24 62.12 63.40 58.11 73.03 70.30 -
-All 60.16 59.90 58.08 57.91 66.90 67.07 67.42
+Domain Distribution Llama2-70B ChatGPT GPT4 Human Baseline Pos Neg All Pos Neg All Pos Neg All All Natural Properties Head 2.78 6.72 0.44 1.45 5.56 0.0 13.68 43.83 9.23 82.31 Long-tail 2.64 3.49 0.0 1.08 5.29 0.0 10.82 39.9 7.21 82.45 $\Delta$100% - - -0.0% - - -21.89% 0.17% Temporal Head 6.58 2.6 0.46 10.72 38.28 4.13 68.3 43.95 36.91 84.69 Long-tail 7.43 3.07 0.0 9.05 32.15 2.26 60.58 38.93 28.11 83.20 $\Delta$100% - - -45.28% - - -23.84% -1.76% Outcomes and Effects Head 7.62 7.93 1.22 19.92 29.27 6.1 57.32 55.18 41.16 83.83 Long-tail 8.5 9.97 0.59 19.65 26.1 2.64 55.72 46.63 33.43 85.13 $\Delta$51.64% - - -56.72% - - -18.78% 1.56% Locational Head 8.57 7.14 0.0 17.14 10.0 5.71 18.57 18.57 2.86 75.71 Long-tail 11.24 8.99 3.37 15.73 4.49 1.12 37.08 8.99 3.37 67.42 $\Delta$ 0.0% - - -80.39% - - 17.83% -10.95% Total Head 5.08 5.28 0.56 8.21 20.67 2.62 39.49 44.87 23.64 83.12 Long-tail 5.69 4.67 0.27 7.76 17.86 1.28 36.58 39.34 18.66 82.44 $\Delta$51.79% - - -51.15% - - -21.07% -0.82%
 Table 4: Performance on the entailment classification task of three LLMs decreases on the long-tail distribution compared to the head distribution, while human performance does not.
-LLMs also have a template bias towards negative templates – when the target token is “no”, “false” and “wrong”.
-4 Can LLM Reason Effectively in the Long-tail Distribution?
-We used all human evaluated knowledge statements that are labeled correct or factually incorrect in LINT, with 1,925 statements in head distribution and 1,856 statements in long-tail distribution, and created a simple entailment classification task to motivate usage of our LINT dataset.
-We evaluate three LLMs: InstructGPT, ChatGPT and GPT4, using both zero-shot and Chain-of-Thought (COT) (Wei et al., 2022) prompting.
-We use data labeled as both factually correct and conforming data type as positive samples and use data labeled as factually incorrect during human evaluation as negative samples.
-For all positive samples, entailment forms between the premise and conclusion, while for negative samples, non-entailment forms.
-Given that LINT has an 84% accuracy, we have more positive samples than negative samples.
-In order to examine whether data distribution affects reasoning behavior in LLMs, we convert each statement into 13 probing questions, where each probing template corresponds to either a positive label (“Yes”, “True”, or “Right”) or a negative label (“No”, “False”, or “Wrong”).
-The probing templates are summarized in Appendix E.1.
+4 LLMs’ (Lack of) Generalization in the Long-tail Distribution
+Using data from LINT, we evaluate LLM generalization through an entailment classification task on inferential knowledge in the long-tail distribution.
+We use all human evaluated knowledge statements except for those with incorrect data types in LINT, with 1,925 statements in head distribution and 1,856 statements in long-tail distribution.
+Statements rated as factually correct has entailment between premise and conclusion, and statements rated as factually incorrect has contradiction between premise and conclusion.
+In order to prevent LLMs’ template biases from misleading the evaluation, we convert each statement into 13 question templates, where each question templates corresponds to a positive label (“Yes”, “True”, or “Right”) or a negative label (“No”, “False”, or “Wrong”).
+The question templates are summarized in Appendix G.1.
+We consider a model answering accurately about one statement only if the model answers all question templates correctly.
+We evaluate three LLMs: llama2-70B, ChatGPT and GPT4.
+In order to enforce the model to predict the target token sets and minimize format noncompliance, we use Chain-of-Thought (CoT) (Wei et al., 2022) prompting that includes 2 in-context examples with randomly shuffled orders of positive label and negative label.
+For each domain, we report aggregated (All) performance of each model as well as human baseline performance in Table 4.
+We also include performance on positive labels only and negative labels only.
+We also mark relative performance drop $\delta=\frac{t-h}{h}$, where $h$ and $t$ are head and tail distribution aggregated performance.
 We obtain human performance on the same set of statements.
@@
 The workers are asked to use search engines to verify their answers.
-See AMT templates in Appendix G.2.
-For each domain, we report aggregated and template-wise performance of InstructGPT, ChatGPT and GPT4 as well as human baseline performance in Table 4.
-Random guessing performance is 54% instead of 50% because of the 13 probing templates.
+See AMT templates in Appendix I.2.
+We make the following observations on LLM generalization in long-tail NLI.
 Performance drops in the long-tail distribution.
-The performance drop of GPT4 is as high as 4% in zero-shot setting and as high as 5% in COT setting.
-Moreover, the long-tail performance of models have a marked gap compared to human baselines, being as high as 19% in the “Natural Properties” domain.
-The positive template performance experience a “long-tail rise” for all models and most domains.
-For all models, the performance on positive templates is significantly lower than on negative templates.
-Besides “Natural Properties” domain, we witness an increase in performance in the long-tail distribution on the positive templates, contrary to our observation on the negative templates.
-We examine the rationale the models generate during COT in Table 12 and find that the models tend to avoid drawing a “definite conclusion”.
-We hypothesize that this phenomenon may be a result of model adjustments made with Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022), and these model adjustments may have a weaker effect on the long-tail distribution due to the knowledge being less frequent during training.
-This hypothesis can be further supported by the significant increase in positive template performance after COT prompting, where COT weakens the model’s original distribution over the answer tokens in the in-context examples.
-We leave the exact explanation to future work.
+All models exhibit a large relative drop in performance in the long-tail distribution.
+The most competitive model, GPT4, has a 21% overall drop from head to long-tail distribution, while other models exhibit a even larger drop.
 Human Performance does not drop for long-tail distribution.
 Performance drop in the long-tail distribution does not occur to humans for 3 out of 4 domains.
-It is expected because humans can verify their knowledge using search engines, so infrequent knowledge does not challenge humans as much as models.
+It is expected because humans can verify their knowledge using search engines, so infrequent knowledge does not challenge humans as much as models (discussion in Appendix A).
 The exception with the locational domain may be due to some relations being less available online(eg.
 banned_in(Food, Country)).
+Brittleness towards question templates.
+The huge gap between model and human baseline performance indicate that LLMs cannot reliably reason on the same statement when question templates change.
+We find that model performance between positive and negative labels can be very different for certain domains, indicating that models are miscalibrated for positive and negative answer tokens.
+Although this phenomenon is not entirely due to the shift of distribution of the knowledge statements, it is caused by model’s unfamiliarity of certain question templates.
+For example, we find that template 5 and template 12 are the same question (“Premise: … Conclusion: … Does the Premise entail the Conclusion?") with opposite answers (“Yes" and “No"), but all models’ performance on template 5 is significantly lower than that on template 12.
+This suggests that models are more likely to create false negatives in such context, another evidence of performance drop due to long-tail distribution.
+In summary, our analysis shows that while human’s inferential reasoning is not affected by their familiarity to the data (provided that they know the entities involved), model’s inferential reasoning ability drops over long-tail knowledge.
+Our result highlights the importance and effectiveness of long-tail evaluation for model generalization, and our dataset LINT can be used as a useful resource for testing inferential generalization of LLMs.
 5 Related Work
-The community has begun to realize the importance of testing language models’ abilities in the long-tail distribution (Godbole and Jia, 2022).
-Recent works reveal that the performance of LLMs is affected by input data probability.
+Works on model generalization analysis have focused on generating adversarial examples for model evaluation (Zhang and Li, 2019; Ziegler et al., 2022; Perez et al., 2022; Casper et al., 2023), flagging abnormal inputs that are likely to trigger bad behavior.
+Recently, the community has realized the importance of testing language models’ abilities in the long-tail distribution (Godbole and Jia, 2022).
+Works reveal that LLM performance is affected by input data probability.
 (McCoy et al., 2023; Razeghi et al., 2022), and more works have focused on generating less common data for probing LLMs.
@@
 (2022) observe a correlation between the model performance on math problems and the frequency of numeric and temporal terms from those instances in the pretraining data.
-In addition to probing models’ reasoning ability on less common data, recent works also propose to test language models’ ability to generate less common dataChen et al.
+In addition to probing models on less common data, recent works also test LLMs generating less common dataChen et al.
 (2023) propose a negative knowledge generation task where models generate uncommon knowledge with negation conditioned on constrained keywords.
@@
 (2023) introduce the “less likely brainstorming” task that asks a model to generate outputs that humans think are relevant but less likely to happen.
-Generating uncommon data is not only challenging for language models, but also challenging for humans because of our cognitive biasTversky and Kahneman (1974) observe that humans are prone to more systematic errors when facing uncertain events, and Tversky and Kahneman (1973) reveal that humans tend to evaluate the frequency of classes or the probability of events by availability, i.e., by the ease with which relevant instances come to mind.
+Generating uncommon data is challenging not only for LLMs, but also for humans because of our cognitive biasTversky and Kahneman (1974) observe that humans are prone to more systematic errors when facing uncertain events, and Tversky and Kahneman (1973) reveal that humans tend to evaluate the frequency of classes or the probability of events by availability, i.e., by the ease with which relevant instances come to mind.
 These traits make it difficult for humans to come up with novel associations (Kray et al., 2006), a crucial ability to create data in the long-tail distribution.
-6 Conclusion and Future Work
-This paper is situated in the problem of generating challenging evaluation data for LLMs in the long-tail distribution.
-We proposed a long-tail data generation framework, LINK, which grounds the statements in symbolic rule templates.
-We showed that zero-shot long-tail knowledge generation is still hard for the SOTA LLMs, and our approach LINK is effective for reaching the long-tail distribution while preserving factual correctness.
-We also showed that LLM’s reasoning capabilities are especially affected in the long-tail regime.
-Using LINK, we created a dataset, LINT, which contains 417 symbolic rules and 54K long-tail knowledge statements.
-LINT is a useful resource for curating challenging evaluation data for LLMs.
-Using a simple entailment classification task, we showed that the long-tail data reveals more flaws of LLM’s reasoning capabilities.
-LINK can facilitate future work on data augmentation.
-One can take natural language statements from existing benchmarks (or new benchmark proposals) and turn them into symbolic rules, from which we can generate long-tail data using LINK.
-Given such long-tail data are less frequently seen during training, it is also useful for mitigate benchmark leakage issues during evaluation.
+6 Conclusion
+Using NLI as a case study, we illustrated the significant potential of long-tail data in uncovering the generalization limitations of LLMs.
+We introduced the first systematic framework designed to generate inferential data within the long-tail distribution, and then demonstrated a noteworthy performance drop of LLMs in the long-tail examples.
+Our work initiates a new line of research focused on long-tail data discovery and generation, urging the research community to adopt long-tail evaluation in the development of LLMs.
 Checklist
@@
 Using LINK on mal-intention-ed rules or searching for toxic and harmful values is a potential threat, but the authors strongly condemn doing so.
-Appendix A Symbolic Rule Creation
-Following the criteria mentioned in § 2.2, we curated 417 symbolic rules from the following sources:
+Appendix A More Discussions
+Our main experiments on entailment classification task assume that the LLMs have the specific knowledge necessary to answer the questions.
+While this is a fair implication, given that their training data are known to contain essentially anything available on the web, one may also worry that the model’s handling of such knowledge is impacted by the memorization of this knowledge.
+As knowledge gets more long-tail , it means it appears in the training data less and is thus harder to memorize.
+Given the imperfect memorization of current LLMs, this may impact their performance on such knowledge, and our current experiments suggest that.
+To verify this, for future experiments one can count the frequency of this long-tail knowledge in the training data to measure how imperfect memorization makes certain knowledge long-tail .
+Appendix B Symbolic Rule Creation
+Following the criteria mentioned in § 2.2, we curated 417 symbolic rules using the following steps:
   1. 1.
-    Expansion from an abstract rule.
-We prompt GPT4 to obtain concrete data types for the abstract rule to create multiple concrete rules.
-Table 5 shows an example of one abstract rule, related prompts for expansion and one concrete rule.
-We obtained 201 rules using this method.
+    Select Compatibility (conclusion will be positive) or Mutual Exclusivity (conclusion will be negative).
   2. 2.
-    Direct generation from concrete data type pairs.
-We prompt GPT4 to generate multiple predicates defining the relationship of the data types as the conclusion, and then generate a premise.
-We obtained 216 rules using this method.
-Table 5: An example meta-rule and its corresponding prompts for expansion.
-Meta-Rule
-is_of_age(Person A, Age X) & requires_a_ minimal_age_of(Object B, Age Y) & is_smaller_than(Age X, Age Y) $\to$ cannot_operate(Person A, Object B)
-Prompt for Data Type Expansion
-In rule “requires_a_minimal_age_of(Object B, Age Y) & cannot_operate(Person A, Object B)", B is a variable representing an object.
+    Defining Constraints.
+Our constraints can be categorized as temporal, locational, natural properties, and desirable/undesirable outcomes and effects.
+Temporal constraints refer to time period or age, locational constraints refer to geographic location such as countries or cities as well as climates (tropical or polar, etc), natural properties refer to physical properties of objects such as temperature, size, density, speed, and outcomes and effects include allergies, cure of disease, etc.
+  3. 3.
+    Selecting argument types. We either use Person-Object or Object-Object as key arguments.
+  4. 4.
+    Defining interactions between arguments.
+Person-Object interactions include “using”, “operating”, “buying”, “consuming”, and Object-Object are more type-specific (eg.
+“scratching” when the constraint is “hardness” in natural property).
+  5. 5.
+    Expanding data types for Person or Object. We prompt InstructGPT to generate more specific data types under Person (eg. Historical Figure) or Object (eg. Vehicle, Tool).
+  6. 6.
+    Optimize the verb describing the interaction. We prompt InstructGPT to generate a more accurate verb for the expanded data types.
+Table 5 shows how an example symbolic rule is constructed.
+Table 5: Illustration of our process for creating an example symbolic rule.
+Constraint Temporal (Age)
+Arguments Person-Object
+Constraint Predicate is_of_age(Person A, Age X) & requires_a_ minimal_age_of(Object B, Age Y) & is_smaller_than(Age X, Age Y)
+Principle Mutual Exclusivity
+Interaction cannot_operate(Person A, Object B)
+Prompt for Data Type Expansion In rule “requires_a_minimal_age_of(Object B, Age Y) & cannot_operate(Person A, Object B)", B is a variable representing an object.
 List 10 subcategories of object that B could be that also make the rule true.
-Expanded Data Types
-Vehicle, Machinery, Alcohol, Firearm, Tattoo Equipment, Tobacco Product
-Prompt for Verb Optimization
-cannot_operate(Person A, Object B) is equal to [mask](Person A, Vehicle B). Write the best predicate that could fit in [mask] token.
-Expanded Rule
-is_of_age(Person A, Age X) & requires_a_ minimal_operating_age_of(Object B, Age Y) & is_smaller_than(Age X, Age Y) $\to$ cannot_drive(Person A, Vehicle B)
-Appendix B Critic Model
+Expanded Data Types Vehicle, Machinery, Alcohol, Firearm, Tattoo Equipment, Tobacco Product
+Prompt for Verb Optimization cannot_operate(Person A, Object B) is equal to [mask](Person A, Vehicle B). Write the best predicate that could fit in [mask] token.
+Expanded Rule Example is_of_age(Person A, Age X) & requires_a_ minimal_operating_age_of(Object B, Age Y) & is_smaller_than(Age X, Age Y) $\to$ cannot_drive(Person A, Vehicle B)
+Appendix C Critic Model
 We find that while the critic model usually verifies data type conformity with high accuracy, it often creates false negatives when verifying factual correctness.
@@
 Table 6: The error rate of each question in human verification. Most errors occur on factual correctness.
-Appendix C Ablation Studies on LINK
-C.1 Hyperparameters on Knowledge Model
+Appendix D Ablation Studies on LINK
+D.1 Hyperparameters on Knowledge Model
 When constructing LINT, we used InstructGPT as the knowledge model with temperature=0.7 and top_p=1.
@@
 Figure 5: All three temperature settings of LINK can successfully generate knowledge statements in the long-tail  distribution, except for when temperature=0.5 in one of the six sampled rules.
-C.2 Effect of Critic on LINK
+D.2 Effect of Critic on LINK
 To investigate the effectiveness of the critic, we provide an ablation study on a few sampled rules by removing the critic in LINK.
@@
 However, the performance drop is less significant in the head distribution Table 9.
-Besides, if we replace the reranker with a random sampling method, the generated statements cannot lie in the long-tail distribution (which will be further explained in § C.3) and have higher correctness without the critic.
+Besides, if we replace the reranker with a random sampling method, the generated statements cannot lie in the long-tail distribution (which will be further explained in § D.3) and have higher correctness without the critic.
 It indicates that it is harder for models to generate correct statements from the long-tail distribution than the head distribution without LINK.
@@
     w/o critic+reranker 88.17 85.33 75.58
-C.3 Effect of Reranker on LINK
+D.3 Effect of Reranker on LINK
 To investigate the effectiveness of the reranker, we provide an ablation study on a few sampled rules by replacing the reranker step with a random sampling method.
@@
 Using reranker is essential for separating the head and long-tail distribution.
-C.4 Ineffectiveness of Post-hoc Reranking for LLM generated knowledge.
+D.4 Ineffectiveness of Post-hoc Reranking for LLM generated knowledge.
 To further highlight the importance of performing step-wise reranking in LINK, we confirm that applying a post-hoc reranker on the GPT4 generations from instructions does not have the same effect as LINK.
@@
 It demonstrates that maneuvering the distribution during the searching process is necessary and more effective than post-hoc filtering.
-C.5 Applying GPT4 as the knowledge model
+D.5 Applying GPT4 as the knowledge model
 Table 10 shows the generation quality of GPT4 using baseline prompting method, LINK and LINK with GPT4 as the knowledge model over 6 sample rules.
@@
 Figure 8: LINK using GPT4 creates statements that fall in a roughly similar long-tail distribution as the original LINK with InstructGPT.
-Appendix D Addendum on Distribution
+Appendix E Addendum on Distribution
 (a) rule0
@@
 While LINK’s long-tail generations fall into a lower probability distribution than those of GPT4, GPT4’s “long-tail distribution" overlaps with the head distribution, indicating that these generations are not truly long-tail.
-D.1 Additional distribution plots for symbolic rules
+E.1 Additional distribution plots for symbolic rules
 As an extension on § 3.2, we show the distribution of statements sampled by LINK, ChatGPT and GPT4 from 6 symbolic rules on InstructGPT in Figure 9.
-D.2 Distribution Comparison of Different Models
+E.2 Distribution Comparison of Different Models
 In this section, we show that the long-tail distribution of different language models overlap, and that this evidence supports our assumption that a universal natural language distribution exists; subsequently, the long-tail distribution of a language model can be used to approximate the long-tail distribution of other language models.
@@
 Figure 12: An illustration of the distribution comparison between llama2-70B and InstructGPT of generated statements by LINK.
-Appendix E Entailment Classification Probing
-E.1 Probing Template
+Appendix F Long-tail Prompts for LLMs
+We tried 10 prompts when prompting LLMs to generate knowledge statements in the long-tail distribution directly with instructions.
+Table 11 shows the 10 prompts which are appended to the original instruction.
+Table 11: An illustration of 10 prompts that instruct LLMs to generate knowledge statements in the long-tail distribution.
+Prompt
+1 Use less frequent terms of A and B and Z.
+2 Use terms of A and B and Z that are less common.
+3 Use terms with lower frequency for A and B and Z.
+4 Use terms of A and B and Z that have lower probability in language model distribution.
+5 Use less frequent words of A and B and Z.
+6 Use words of A and B and Z that are less common.
+7 Use words with lower frequency for A and B and Z.
+8 Use less frequent entities of A and B and Z.
+9 Use entities of A and B and Z that are less common.
+10 Use entities with lower frequency for A and B and Z.
+Figure 13: LINT has a higher $\delta$ than GPT4 baselines with 10 different prompts.
+Following Section 3.2, Figure 13 shows the $\delta$ value for GPT4 baselines with these prompts and LINT.
+These prompts have a similar effect on the distribution of generated statements on most rules and LINT consistently has a higher $\delta$, indicating that no matter the prompt we use, LLMs cannot directly generate long-tail statements by following instructions.
+Though these prompts have different $\delta$ values on some rules, no one can consistently achieve a higher $\delta$ among them.
+Appendix G Entailment Classification Probing
+G.1 Probing template
 Table 12 shows templates we used for the entailment classification task.
@@
 13 Premise: premise. Conclusion: conclusion. Does premise contradict the conclusion? No
-E.2 Rationale analysis
+G.2 Accuracy averaged over templates
+One main concern of not using accuracy of our main metric is because accuracy is not favorable for label imbalance.
+Because we have 13 templates where for each knowledge statement we will have an uneven number of positive or negative answers, the negative label rate in total is around 51.58%.
+Therefore, we decide to consider the template biases and deem a model correct only if it answers all templates of a statement correctly, as our goal is to test the true knowledge of models.
+When evaluating the accuracy across all templates, the model’s performance also drops over long-tail knowledge.
+Table 13 shows the accuracy across all templates of GPT4.
+Even if LLMs are brittle to templates, they exhibit a performance drop in the long-tail distribution among all domains.
+Head Long-tail
+Natural Properties 67.07 63.40
+Temporal 83.14 78.97
+Outcomes and Effects 81.79 80.19
+Locational 68.90 67.07
+Table 13: Accuracy over all templates, GPT4
+G.3 Model and human error analysis
+Table 14 shows probing examples in which models and humans make error.
+For those from the head distribution, the entailment can be easy if humans can use search engines.
+For example, humans can search the periods of “The Paleolithic Era” and “Lion Gate of Mycenae”, and answer easily.
+Thus the human errors in the head distribution may be due to carelessness.
+For those containing long-tail knowledge, even with search engines, it is not so easy to infer the answer for human annotators.
+It is also likely that models do not have such knowledge either.
+Table 14: Examples that human and model make mistakes.
+Statement: Yes or no: if Plant X vanished in Paleolithic Era, Plant X cannot surround Lion Gate of Mycenae. (Head)
+True label: Yes
+Human answer: No
+Model answer (GPT4 + COT): Yes.
+The Paleolithic Era ended around 10,000 BC and the Lion Gate of Mycenae was built around 1250 BC.
+If PlantX vanished in the Paleolithic Era, it could not have surrounded the Lion Gate of Mycenae which was built much later.
+Statement: Yes or no: if Bag X has trouble containing Pandeiro, Dhak cannot fit in Bag X. (Long-tail )
+True label: Yes
+Human answer: No
+Model answer (GPT4 + COT): No.
+The size of the Pandeiro does not directly relate to the size of the Dhak.
+They are different objects and may have different shapes and sizes.
+Therefore, just because BagX has trouble containing a Pandeiro doesn’t necessarily mean that a Dhak cannot fit in BagX.
+G.4 Rationale analysis
 As mentioned in § 4, we examine the rationale the model generated during COT in the entailment classification task and found that the models tend to avoid drawing a “definite conclusion”.
@@
 Table 15: A rationale generated by GPT4 with COT shows that even though the model is able to cite the relevant information, it inclines to answer “No" because it cannot “definitively conclude".
-Input
-Premise: PersonX is of age 17 years 8 months. Conclusion: PersonX cannot consume Bourbon. Does premise entail conclusion?
-Label
-Yes
-Answer
-No
-Rationale
-In many countries, the legal drinking age is 18 or 21. If we do not know the specific country PersonX is in, we cannot definitively conclude that they cannot consume Bourbon.
-Appendix F Domain-wise statistics of LINT
+Input Premise: PersonX is of age 17 years 8 months. Conclusion: PersonX cannot consume Bourbon. Does premise entail conclusion?
+Label Yes
+Answer No
+Rationale In many countries, the legal drinking age is 18 or 21. If we do not know the specific country PersonX is in, we cannot definitively conclude that they cannot consume Bourbon.
+Appendix H Domain-wise statistics of LINT
 Table 16 shows the detailed statistics of LINT.
@@
 Temporal 81 15,143 15,317
-Capability and Advice 132 14,966 15,010
+Outcomes and Effects 132 14,966 15,010
 Natural Properties 139 16,788 16,669
@@
 Total 417 54,220 54,366
-F.1 Domain-wise human evaluation
+H.1 Domain-wise human evaluation
 As mentioned in § 3.3, we uniformly sample 4,000 statements from LINT for human evaluation.
@@
 Temporal 90.18 85.27 77.38
-Capability and Advice 94.97 80.73 75.98
+Outcomes and Effects 94.97 80.73 75.98
 Natural Properties 96.61 96.61 93.81
 Locational 98.88 70.79 70.79
-F.2 Rule definitions
+H.2 Rule definitions
 Table 18 shows the definitions of the 6 sampled rules.
 Table 18: Rule definitions of six sampled rules.
-Rule0
-lived_in(Person P, Geographic Location A) & lived_during(Person P, Historical Time Period D) & existed_during(Geographic Location A, Historical Time Period D) & was_invented_in(Product or Technology C, Year Y) & is_more_than_a_century_earlier_ than(Historical Time Period D, Year Y) $\to$ is_not_able_to_use(Person P, Product or Technology C)
-Rule32
-has_trouble_lifting(Person X, Name of Appliance B) & is_heavier_than(Object A, Name of Appliance B) $\to$ cannot_lift(Person X, Object A)
-Rule46
-is_allergic_to(Person A, Substance X) & includes(Name of Cosmetics B, Substance X) $\to$ cannot_use(Person A, Name of Cosmetics B)
-Rule88
-died_in(Historical Figure A, Historical Time Period X) & was_created_during(Artifact B, Historical Time Period Y) & is_earlier_than( Historical Time Period X, Historical Time Period Y) $\to$ cannot_create(Historical Figure A, Artifact B)
-Rule112
-has_trouble_containing(Room B, Furniture C) & is_larger_than(Furniture A, Furniture C) $\to$ cannot_fit_in(Furniture A, Room B)
-Rule122
-has_trouble_containing(Trunk B, Furniture C) & is_larger_than(Furniture A, Furniture C) $\to$ cannot_fit_in(Furniture A, Trunk B)
-Appendix G Amazon Mechanic Turk
-G.1 Recruiting Workers
+Rule0 lived_in(Person P, Geographic Location A) & lived_during(Person P, Historical Time Period D) & existed_during(Geographic Location A, Historical Time Period D) & was_invented_in(Product or Technology C, Year Y) & is_more_than_a_century_earlier_ than(Historical Time Period D, Year Y) $\to$ is_not_able_to_use(Person P, Product or Technology C)
+Rule32 has_trouble_lifting(Person X, Name of Appliance B) & is_heavier_than(Object A, Name of Appliance B) $\to$ cannot_lift(Person X, Object A)
+Rule46 is_allergic_to(Person A, Substance X) & includes(Name of Cosmetics B, Substance X) $\to$ cannot_use(Person A, Name of Cosmetics B)
+Rule88 died_in(Historical Figure A, Historical Time Period X) & was_created_during(Artifact B, Historical Time Period Y) & is_earlier_than( Historical Time Period X, Historical Time Period Y) $\to$ cannot_create(Historical Figure A, Artifact B)
+Rule112 has_trouble_containing(Room B, Furniture C) & is_larger_than(Furniture A, Furniture C) $\to$ cannot_fit_in(Furniture A, Room B)
+Rule122 has_trouble_containing(Trunk B, Furniture C) & is_larger_than(Furniture A, Furniture C) $\to$ cannot_fit_in(Furniture A, Trunk B)
+Appendix I Amazon Mechanic Turk
+I.1 Recruiting Workers
 We recruit workers from all English-speaking countries (US, UK, New Zealand, Australia, Canada), although AMT workers are mostly US-based.
@@
 In the qualification task, all workers will be presented with three manually selected statements, which are clear and representative.
-Each statement has five related questions as described in Appendix B.
+Each statement has five related questions as described in Appendix C.
 Only workers who answer all the questions correctly will be recruited.
@@
 We paid the workers \$0.47 per annotation for evaluating the quality of generations and \$0.11 per annotation for the entailment classification task, to match \$15 per hour based on their working time.
-G.2 Templates
+I.2 Templates
 Figure 14 and Figure 15 show the template we use for the evaluation of generation quality and the entailment classification task.
-G.3 Agreement statistics
+I.3 Agreement statistics
 Table 19 shows the agreement of annotations in the evaluation task.
@@
 Overall 65.64 74.93 83.39
-G.4 Failure Case Examples
+I.4 Failure Case Examples
 We analyze some failure cases that are labeled as incorrect in the human evaluation. Table 20 presents some examples.
 Table 20: Examples that are labeled as incorrect during human evaluation. Note that the reasons are analyzed by the authors instead of annotators.
-Rule 172
-Locational
-Rule: is_located_in(Person A, Location X) & is_forbidden_in(Food Item B, Location X) $\to$ cannot_eat(Person A, Food Item B) Premise: Person X is located in Houston Conclusion: Person X cannot eat Chocolate Is Houston a location?
+Rule 172 Locational Rule: is_located_in(Person A, Location X) & is_forbidden_in(Food Item B, Location X) $\to$ cannot_eat(Person A, Food Item B) Premise: Person X is located in Houston Conclusion: Person X cannot eat Chocolate Is Houston a location?
 Annotation: Yes Is Chocolate a food item?
@@
 Chocolate is not actually forbidden in Houston, so People in Houston can eat chocolate.
-Rule 371
-Capability and Advice
-Rule: can_treat(Drug B, Name of Disease X) & has(Person A, Name of Disease X) $\to$ should_take(Person A, Drug B) Premise: Person X has Hepatitis Conclusion: Person X should take Sofosbuvir Is Hepatitis a name of disease?
+Rule 371 Capability and Advice Rule: can_treat(Drug B, Name of Disease X) & has(Person A, Name of Disease X) $\to$ should_take(Person A, Drug B) Premise: Person X has Hepatitis Conclusion: Person X should take Sofosbuvir Is Hepatitis a name of disease?
 Annotation: Yes Is Sofosbuvir a drug?
@@
 For other types of hepatitis, different medications or treatments may be necessary.
-Rule 274
-Temporal
-Rule: vanished_in(Plant A, Historical Time Period X) & was_invented_in(Weapon B, Historical Time Period Y) & is_earlier_than(Historical Time Period X, Historical Time Period Y) $\to$ cannot_be_used_to_conceal(Plant A, Weapon B) Premise: Plant X vanished in Mongol Conclusion: Plant X cannot be used to conceal M92 Zolja Is Mongol a historical time period?
+Rule 274 Temporal Rule: vanished_in(Plant A, Historical Time Period X) & was_invented_in(Weapon B, Historical Time Period Y) & is_earlier_than(Historical Time Period X, Historical Time Period Y) $\to$ cannot_be_used_to_conceal(Plant A, Weapon B) Premise: Plant X vanished in Mongol Conclusion: Plant X cannot be used to conceal M92 Zolja Is Mongol a historical time period?
 Annotation: No Is M92 Zolja a weapon?
@@
 The Mongol Empire may refer to a period of the 13th and 14th centuries, but Mongol cannot.
-Rule 204
-Natural Properties
-Rule: has_trouble_containing(Drawer B, Tool C) & is_larger_than (Tool A, Tool C) $\to$ cannot_be_placed_in(Tool A, Drawer B) Premise: Drawer X has trouble containing Scroll saw Conclusion: Car cannot be placed in Drawer X Is Scroll saw a Tool?
+Rule 204 Natural Properties Rule: has_trouble_containing(Drawer B, Tool C) & is_larger_than (Tool A, Tool C) $\to$ cannot_be_placed_in(Tool A, Drawer B) Premise: Drawer X has trouble containing Scroll saw Conclusion: Car cannot be placed in Drawer X Is Scroll saw a Tool?
 Annotation: Yes Is Car a Tool?
@@
 Car is a vehicle instead of a tool.
-Refer to caption Figure 13: AMT template for the evaluation of generation quality. Refer to caption Figure 14: AMT template for the human baseline of the entailment classification task.
+Figure 14: AMT template for the evaluation of generation quality.
+Figure 15: AMT template for the human baseline of the entailment classification task.